### v1

# Synthetic Data Generation in Cybersecurity: A Comparative Analysis

Abstract
Synthetic data generation faces significant challenges in accurately replicating real data, particularly with tabular data, where achieving high fidelity and utility is critical. While numerous methods have been developed, the most effective approach for creating high-quality synthetic data for network traffic security remains to be seen. This study conducts a comprehensive comparative analysis of non-AI, conventional AI, and generative AI techniques for synthetic tabular data generation using two widely recognized cybersecurity datasets: NSL-KDD and CICIDS-2017. Particular emphasis was placed on prominent GAN models for tabular data generation, including CTGAN, CopulaGAN, GANBLR++, and CastGAN. The results indicate that GAN-based methods, particularly CTGAN and CopulaGAN, outperform non-AI and conventional AI approaches in terms of fidelity and utility. To the best of our knowledge, this research contributes to the field by offering the first comparative evaluation of these methods specifically for cybersecurity network traffic data, filling a critical gap in the literature. It also introduces mutual information for feature selection, further enhancing the quality of the generated synthetic data. These findings provide valuable guidance for researchers seeking the most suitable synthetic data generation method in cybersecurity applications.
Index Terms:
Generative Adversarial Networks (GANs), Synthetic tabular Data, Cybersecurity, NSL-KDD, CICIDS, Mutual information, Feature selection
I Introduction
In the rapidly evolving field of cybersecurity, IDS plays a crucial role in identifying and mitigating threats within network environments. These systems rely heavily on high-quality data to detect patterns indicative of malicious activity [1]. However, acquiring real-world cybersecurity data poses significant challenges, primarily due to privacy concerns and time constraints [2]. Researchers have turned to synthetic data generation as a viable alternative to address this. Synthetic data, when adequately generated, can mimic the characteristics of real-world data while avoiding the associated ethical and legal issues, making it a valuable resource for training and evaluating IDS [3].
Despite the growing interest in synthetic data generation, producing synthetic datasets that maintain the utility, fidelity, and data integrity required for practical cybersecurity applications remains a significant challenge [4]. Utility refers to the ability of synthetic data to support the same analytical tasks as real data [5]. At the same time, fidelity ensures that the synthetic data accurately reflects the statistical properties of the original data[6]. Data integrity, on the other hand, focuses on the accuracy and reliability of data, ensuring that it remains unaltered and consistent [7, 8]. Follow these criteria to avoid synthetic data that is either unrealistic or similar to real data, rendering it less valuable or risky.
GANs have shown great potential for generating synthetic tabular data for cybersecurity purposes. Various GAN architectures like VanillaGAN[9], WGAN[10], and WGAN-GP[11], as well as models specifically designed for tabular data generation such as CTGAN[12], CopulaGAN [13], and GANBLR [14], have been successfully used to create synthetic IDS data [15]. However, the effectiveness of GANs for generating tabular data specific to cybersecurity, such as network flow data used in IDS, has yet to be thoroughly benchmarked against other tabular synthetic data generation methods: Artificial intelligence (AI) based and non-AI based.
I-A Synthetic data for cybersecurity
  1. 1.
    Data privacy: Sharing real-world cybersecurity data involves a risk of sharing sensitive information, which hinders collaboration and research [16]. Thus, the synthetic data resolves this issue, and valuable data can be shared without compromising privacy [17].
  2. 2.
    Data scarcity: The real-world cyber security data is often imbalanced, with normal traffic making up a higher percentage than attacks or malicious traffic. This causes a problem of data scarcity. Synthetic data in this scenario can help balance both instances and improve cyber security tasks, such as efficient, unbiased intrusion detection [1].
  3. 3.
    Diversity of scenarios: Cybersecurity threats are becoming more dynamic and diverse, but real cybersecurity datasets often need more diversity [18]. In this scenario, synthetic data helps simulate various uncommon attacks and allows for more comprehensive testing of cybersecurity tools [19].
  4. 4.
    Enhancing AI models: When training cyber security models such as Intrusion detection systems (IDS), a large, high-quality dataset is essential for effective detection. Therefore, synthetic data plays a crucial role in training these models on a diverse and extensive dataset to enhance the robustness of AI models [20, 21].
I-B Synthetic data challenges in cybersecurity
  1. 1.
    Data realism: One of the primary challenges in generating synthetic data is ensuring its realism compared to real-world data [22]. The effectiveness of AI models trained on synthetic data depends on how accurately the data represents real cyber threats and network conditions[23]. Many generative models are still in the process of achieving data realism by increasing their fidelity and utility to mimic actual trends [24].
  2. 2.
    Lack of variability: Some generative models are explicitly designed to generate a diverse set of network instances and cyber threats. However, if synthetic data is not generated with sufficient variability, it may miss capturing the full spectrum of possible attack vectors in cybersecurity [25, 26]. For instance, the Wasserstein GAN has solved the problem faced by the vanilla GAN of mode collapse, thus ensuring that the generated outcome is not limited to a few instances [10].
  3. 3.
    Class imbalance: Cybersecurity often suffers from class imbalance, where regular traffic is much larger than cyberattacks [27]. Synthetic data generation methods address this problem by balancing malicious and benign traffic and attacks within the cybersecurity dataset for more robust intrusion detection [28].
  4. 4.
    Privacy concerns: Synthetic data attempts to conceal confidential information from real-world sensitive data to address privacy and security concerns [29]. However, it is not entirely free from these challenges. Ensuring that sensitive information cannot be traced back from the synthetic data is crucial and remains a concern.
  5. 5.
    Standardization: More standardization is needed for generating, validating, and integrating synthetic data into cybersecurity practices. This makes it challenging to ensure the reliability and consistency of the synthetic datasets across different cybersecurity applications [30, 31].
  6. 6.
    Perfomace evaluation: For evaluating synthetic data, it is crucial to note that achieving a high similarity score does not guarantee its real-world applicability. Similarly, obtaining high accuracy in synthetic cybersecurity datasets for any ML/DL model does not ensure compatibility with real-world datasets [24, 32].
I-C Motivation
This research addresses critical challenges in strengthening IDS defenses amidst a rapidly escalating cyber threat landscape. IDS are vital for securing networks by detecting and mitigating intrusions, but a lack of high-quality, real-world data for training and refinement hampers their effectiveness. Obtaining such data is challenging due to privacy concerns, ethical issues, and evolving network behaviors. Synthetic data generation using techniques like GANs offers a promising solution by creating realistic, high-fidelity data that mimics network activities without violating privacy or legal constraints. This research is crucial to improving IDS capabilities, enabling them to adapt to modern cyber adversaries’ sophisticated tactics. This study aims to enhance IDS robustness and reliability by benchmarking different synthetic data generation methods, contributing to more secure cybersecurity frameworks.
This research fills a gap in cybersecurity by offering a comprehensive comparison of synthetic data generation methods for enhancing IDS. While synthetic data’s potential has been recognized, there is limited clarity on which methods—traditional, AI-based, or GAN-based—provide the highest fidelity and utility for IDS. Existing literature has explored various techniques but lacks a focused comparison of their effectiveness in generating network traffic data for IDS. This study bridges the gap by systematically evaluating different synthetic data generation techniques, guiding the development of more effective IDS and contributing to more robust, more adaptable cybersecurity defenses.
I-D Objectives
The primary objective of this study is to rigorously evaluate and compare the performance of various synthetic data generation methods for cybersecurity, focusing on their utility, fidelity, data integrity, and ability to handle class imbalance. To achieve this, the research addresses the following key questions:
  1. 1.
    Which methods produce synthetic data that best preserves the utility for training IDS models?
  2. 2.
    How well do these methods maintain the original data’s statistical properties (fidelity) and integrity?
  3. 3.
    Which techniques effectively balance class distributions in the generated data?
  4. 4.
    Which GAN models are best for producing cyber security network traffic data? These questions guide the comprehensive benchmarking of both statistical and AI-based synthetic data generation techniques.
I-E Contribution
The contributions of this work are summarized below:
  1. 1.
    Comprehensive Comparison of Synthetic Data Generation Methods: This research uniquely compares three prominent synthetic data generation methods—statistical, classical AI, and generative AI—on cybersecurity network traffic datasets (NSLKDD, CICIDS17). This provides the first direct comparison of these methods in the cybersecurity field.
  2. 2.
    Filling the Knowledge Gap in Cybersecurity Data: The study addresses a gap in the literature by identifying the most suitable synthetic data generation method specifically for network traffic datasets, offering clarity in an area where no previous comparative study exists.
  3. 3.
    Practical Guidance for Researchers: By evaluating the performance of these methods, this research offers a clear recommendation on the best approach for generating synthetic cybersecurity data, enabling researchers to make informed decisions without having to test multiple methods.
  4. 4.
    Introduction of Mutual Information for Feature Selection: Using Mutual Information for feature selection in synthetic data generation is an innovative contribution, providing a new perspective that may improve feature selection strategies in future cybersecurity studies.
The source code of this experiment is publicly available on our GitHub Repository  [^1] https://github.com/AdenRajput/Comparative_Analysis.git.
I-F Paper Organization
The remaining sections of this paper are organized as follows:
  • Section II provides an overview of Synthetic Tabular Data Generation Methods, their usage in tabular data applications in cybersecurity, and evaluation metrics.
  • Section III details the prominent standard non-AI and AI-based synthetic tabular data generation methods and their mathematical workflow.
  • Sections IV and V describe the mutual information and evaluation metrics used in this experiment.
  • Section VI discusses the experimental setup, including the datasets and model configurations.
  • Section VII presents our experiments’ results, highlighting each method’s performance across two datasets.
  • Finally, in Sections VIII and IX, we mention the potential areas for future work and development, focusing on critical aspects that require further investigation and improvement, and conclude with key highlights of this study.
II Related Work
II-A Synthetic Tabular Data Generation Methods
Synthetic data is an artificially produced dataset that mimics the statistical features of real-world data without disclosing any sensitive information [3]. This data type spans various forms, such as tabular data, images, and text.
Tabular data, organized in a table format with rows and columns, is widely used for ML and DL modeling. It is also crucial for cybersecurity applications due to its structured nature. It is ideal for representing logs, transaction records, and other important information for detecting and mitigating security threats [16]. For example, cybersecurity datasets often contain tabular data in IDS, fraud detection, and anomaly detection, enabling precise analysis and decision-making [33, 34]. Additionally, tabular data’s clear and organized format allows for effectively applying ML and DL algorithms to enhance security measures [22].
The significance of synthetic data in cybersecurity stems from the challenges associated with accessing high-quality, real-world data in this field. Privacy concerns, regulatory limits, and the sensitive nature of cybersecurity information often restrict the availability of data [35]. By facilitating the generation of realistic datasets free of genuine, sensitive information, synthetic data offers a solution to these challenges [36]. This advancement is critical for pushing forward research, enhancing security models, and executing thorough threat analyses.
II-B Applications in Cybersecurity
Synthetic data finds multiple applications in cybersecurity, improving different facets of security management and threat response:
  1. 1.
    Intrusion Detection Systems (IDS): By supplying varied and representative datasets encompassing a range of attack patterns and behaviors, synthetic data aids in the training and testing of IDS, enhancing their detection accuracy and resilience [37]. Research has shown that employing synthetic data techniques on datasets like UNSW-NB15 and CICIDS-17 has significantly improved the performance of deep learning models used in IDS [38].
  2. 2.
    Anomaly Detection: Synthetic data plays a role in simulating typical and atypical behaviors within networks, thereby facilitating the training of systems to detect deviations indicative of potential security incidents [39]. Synthetic data has been used to preserve privacy and detect anomalies in the KDD dataset. Its effectiveness, measured in supervised, semi-supervised, and unsupervised anomaly detection, showed an accuracy of 99.69%, similar to models trained on original data [40].
  3. 3.
    Security Training and Simulation: The value of synthetic data extends to crafting realistic scenarios for security training exercises and simulations. This enables security teams to hone their responses to varied cyber threats without compromising real data [41]. A deep Convolutional Generative Adversarial Network (DCGAN) model was used to generate synthetic cyberattack data and realistic attack simulations to train deep learning models for better detection. The model with those simulated attacks achieved an accuracy of 99.69% on the KDD dataset and 97.93% on the CICIDS-17 dataset [42].
In discussing the significance and uses of synthetic tabular data in cybersecurity, we will focus on the methods used to generate such data effectively. The later sections will delve into the specifics and their architecture. They will also provide an experimental analysis of the prominent standard (non-AI) and AI-based methods for creating synthetic tabular data for cybersecurity IDS.
II-C Metrics for Evaluation
  1. 1.
    Fidelity: Fidelity measures how well the synthetic data replicates the statistical properties of the original real-world data [43]. Therefore, the single most significant metric for evaluating synthetic tabular data generally involves assessing the data’s underlying distribution [44]. This approach is based on the definition of data synthesis to capture the probability distribution function of the original data, thereby creating synthetic data that mirrors these characteristics [45]. However, the evaluation process can be subjective, based on the data production’s objective. As a result, researchers employ various methods to assess synthetic tabular data, considering its structure markedly differs from that of images.
    Unlike synthetic tabular data, the validation of synthetic image-based output is often visual. If the model yields realistic images, it’s deemed successful. Images comprise spatial information, shape, texture, objects, and color [46]. Nonetheless, synthetic data is confined to images and encompasses tabular data [47], characterized by its complexity, multiple interconnected features, hidden patterns, and context-dependent information [48]. Hence, evaluating synthetic tabular data demands a more structured approach.
  2. 2.
    Utility: Utility measures the usefulness of synthetic data for the same tasks or analyses as the original data [49]. To evaluate the quality of the synthetic data, key performance metrics such as accuracy, F1 score, precision, and recall were compared between models trained and tested on real data versus those involving synthetic data[50, 15, 51]. Specifically, the following scenarios were used to rank performance:
    TRTR (Train and Test on Real Data): Baseline performance using real data. TSTR (Train on Synthetic, Test on Real): Measures how well models trained on synthetic data generalize to real data. TRTS (Train on Real, Test on Synthetic): Assesses the consistency of synthetic data in replicating real data behavior. TSTS (Train and Test on Synthetic Data): Evaluates synthetic data’s internal consistency and utility within a self-contained model [52, 53]. By comparing these four scenarios, the utility of the synthetic data can be ranked, providing insight into how well it replicates the practical use of real data [54, 55, 56].
  3. 3.
    Universal metrices: Recent research has examined synthetic tabular data via diverse statistical and ML methods [57, 58, 59]. Nevertheless, these approaches could be more precise regarding suitable evaluation measures. Hence, there’s been a move to define universal metrics to boost universality and interpretability. One such metric is TabSynDex, offering a score between 0 and 1 to gauge the synthetic data’s closeness to real data [60]. However, this metric still needs to improve in selecting the proper evaluation measures. For example, there needs confirmation that TabSynDex’s five main components adequately represent the synthetic data’s completeness or facilitate more accessible validation of generated tabular data. A recent SCH Yang et al. proposal introduces a structured evaluation methodology focused on comparing the distributions of synthetic and real data [44]. This study is pivotal as it aligns with generative AI’s core principle: replicating the probability distribution of the original dataset.
Therefore, by considering the recent research in the domain, this study employs prominent evaluation metrics such as fidelity (statistical similarity), utility, and class balance in its evaluation section.
III Synthetic Data Generation Techniques
Figure 1: Methods for generating synthetic tabular data
III-A Standard Methods (non-AI)
  1. 1.
    Random Oversampling (ROS) is a widely used technique to address class imbalance, a common issue in datasets, especially in domains like cybersecurity where certain types of data, such as network intrusions, are rare compared to normal traffic. ROS works by duplicating samples from the minority class to balance the class distribution, thereby preventing ML models from becoming biased toward the majority class [61]. This technique is particularly relevant when the non-complex small dataset lacks sufficient examples of minority classes.
    In synthetic tabular data generation, ROS is often employed to artificially increase the number of minority class samples, helping models better learn the characteristics of these rare events. However, its application in cybersecurity data, which often involves complex relationships between variables, may be limited if used alone, as it does not generate new, diverse data points [62].
    The basic procedure of ROS is straightforward:
    1. (a)
      Identify the Minority Class: Determine which class has fewer samples.
    2. (b)
      Duplicate Samples: Randomly select samples from the minority class and duplicate them until the class distribution is balanced.
    3. (c)
      Augmented Dataset: The result is a dataset with a balanced class distribution, where the minority class is equally represented as the majority class
    This process does not alter the duplicated samples, meaning the new data points are exact copies of the original samples.
  2. 2.
    Synthetic Minority Oversampling Technique (SMOTE) is an advanced over-sampling method designed to address the limitations of ROS by generating synthetic examples rather than merely duplicating existing ones. Introduced by Chawla et al. in 2002, SMOTE aims to balance the class distribution in datasets by creating new instances of the minority class, which helps improve the performance of ML models, particularly in scenarios where the class imbalance is significant [63]. Unlike traditional over-sampling methods, SMOTE synthesizes new samples by interpolating between existing minority class samples, leading to more generalized decision boundaries for classifiers [64].
    SMOTE is particularly useful in domains such as cybersecurity, where the occurrence of attacks (minority class) is much rarer than normal behavior (majority class). The technique is well-suited for situations where the minority class instances are spread sparsely across the feature space, and duplicating existing samples (as done in ROS) could lead to overfitting [65, 66].
    The SMOTE algorithm operates through the following steps:
    1. (a)
      Minority Class: For each minority class instance, there is a need to select k-nearest neighbors, e.g., k = 5
    2. (b)
      Synthetic Samples: One of the k-nearest neighbors is randomly chosen to generate a synthetic data point. The synthetic data point is created by interpolating between the feature vector of the original data point and that of the selected neighbor. This interpolation involves calculating the difference between the feature vector of the selected neighbor and the original data point. Then, this difference is multiplied by a random number between 0 and 1, and the resulting value is added to the feature vector of the original data point.
    3. (c)
      Augmented Dataset: The result is a dataset augmented with synthetic samples that exhibit variations within the minority class, thereby mitigating the risk of overfitting and improving the model’s ability to generalize.
  3. 3.
    Adaptive Synthetic Sampling (ADASYN) is an oversampling method similar to ROS and SMOTE. However, it introduces a more adaptive approach to generating synthetic data. It was introduced by He et al. in 2008 and focuses on the harder-to-learn examples from the minority class. ADASYN improves on SMOTE by adapting the sampling procedure based on the learning difficulty of each minority class sample. Data points that are harder to classify receive more synthetic neighbors; the rest receive fewer neighbors. [67].
    ADASYN is particularly useful in domains such as cybersecurity, where the rarity of specific attack types (minority class) can cause models to perform poorly [68]. By adaptively generating synthetic samples based on the distribution of the minority class, ADASYN helps models learn the decision boundary more effectively, leading to better generalization and improved detection of minority class instances [69].
    ADASYN operates through the following steps:
    1. (a)
      Identify Difficult Minority Class: ADASYN computes the local density of each minority class by analyzing how many of its nearest neighbors belong to the majority class. Data points surrounded by majority-class neighbors are considered harder to classify.
    2. (b)
      Weight Assignment: Weights are assigned to each minority instance/data point based on difficulty level. Harder-to-classify points receive a higher weight, implying that more synthetic samples will be generated.
    3. (c)
      Synthetic Samples: For each minority instance, ADASYN interpolates between the instances and one of its nearest minority class neighbors. Synthetic samples are created along the line connecting the two instances.
    4. (d)
      Data Balancing: ADASYN adds the generated samples to the original ones, focusing on the problematic region to reduce the class imbalance and improve the model training and learning.
  4. 4.
    Cluster Centroids is an under-sampling technique designed to address class imbalance by reducing the number of majority class instances. Unlike traditional under-sampling methods, which randomly remove majority class samples, Cluster Centroids utilizes a clustering algorithm to identify representative centroids within the majority class. The technique was popularized to preserve the majority class’s distribution and structure while reducing its size, thereby balancing the dataset without losing critical information [70, 71]. This approach is beneficial in scenarios where the majority class is substantial, and random under-sampling could significantly lose valuable data.
    The concept of cluster centroids is applicable in various domains, including cybersecurity. In highly imbalanced datasets, where normal traffic significantly outweighs the attack data, cluster centroids help by grouping similar attack (minority class) samples into clusters and generating synthetic data points at the centroid of each cluster. This helps represent the attack patterns more effectively and exposes ML models to diverse attacks for robust anomaly detection [72].
    Cluster Centroids operates through the following steps:
    1. (a)
      Minority Class Clustering: Minority class data is grouped into clusters using a clustering algorithm, e.g., K-Means clustering. Each cluster contains similar data points from the minority class.
    2. (b)
      Centroid Calculation: For each cluster, the algorithm computes the centroids, the central points representing the average position of the data points within a particular cluster in the feature space.
    3. (c)
      Synthetic Data: Synthetic samples are generated at or around the center of each minority cluster. This ensures that the generated data is representative of the overall distribution of the minority class.
    4. (d)
      Balanced Dataset: Synthetic data is added to the original data to balance class distributions, improving the AI model’s ability to learn from minority classes for higher efficiency.
    Figure 2: Bayesian Network illustrating some key feature dependencies in the NSL-KDD dataset
  5. 5.
    Gaussian Mixture Model (GMM) is a probabilistic model used to represent the presence of sub-populations within an overall population without requiring that an instance belong exclusively to a single sub-population. GMMs are particularly useful for modeling data that exhibit multiple underlying distributions, assuming that all data points are generated from a mixture of several Gaussian distributions with unknown parameters [73]. Each Gaussian component within a GMM represents a cluster characterized by its mean, variance, and mixing coefficient.
    In the context of synthetic tabular data generation, GMMs approximate the data distribution, making them highly applicable for generating new samples that preserve the statistical properties of the original dataset. This is especially useful in cybersecurity, where modeling the intricate patterns of network traffic or attack vectors requires flexible and accurate representation of complex, multimodal distributions [74]. GMMs provide a robust method for capturing these patterns, allowing for generating synthetic data that closely mimic real-world scenarios [75].
    GMM operates through the following steps:
    1. (a)
      Fitting GMM: GMM fits the original data by assuming it comes from the mixture of multiple Gaussian distributions. It estimates these Gaussian components’ parameters (mean, covariance, and weights) using the Expectation-Maximization (EM) algorithm.
    2. (b)
      Estimate Gaussian Parameters: The model assigns data points to different Gaussian components based on their likelihood and calculates the mean (center), covariance (spread), and mixture weight (probability and skewness) for each Gaussian distribution.
    3. (c)
      Sampling: After training the model, synthetic data is generated by randomly selecting a Gaussian component (based on its weight) and then sampling a point from the selected Gaussian using its parametric values.
    4. (d)
      Data Generation: Repeating the sampling process multiple times generates synthetic data that mimics the structure and underlying distribution of the original data. Thus, the generated data points from all components are combined to form the final synthetic dataset. [76].
Figure 3: Variational Autoencoder Structure
    III-B AI based Methods
  1. 1.
    Baysian Network (BN) are probabilistic graphical models that represent a set of variables and their conditional dependencies via a directed acyclic graph (DAG) [77]. Bayesian Networks are particularly effective for modeling complex relationships in data where uncertainty and causality play critical roles [78, 79]. In figure 2, this BN illustrates that target is a central feature, influenced by several other variables, including service_http, service_private, flag_SF, logged_in, and flad_S0. This suggests that target plays a crucial role in determining the system’s behavior, as it is obvious by its binary data type of normal or attack. The BN also shows key conditional dependencies, such as flag_SF being dependent on logged_in and service_domain_u, while dst_host_serror_rate influences flad_S0. Some features, like service_private and service_http, directly impact target but do not further influence other variables. Overall, this network emphasizes the significant joint dependencies between these system features, highlighting target as a critical node and capturing the complex relationships between the various network-related parameters.
    In synthetic data generation, BNs generate data that preserves the inherent dependencies and conditional probabilities observed in the original dataset. This is particularly valuable in cybersecurity, where understanding the intricate dependencies between various network events, user behaviors, or attack patterns is essential for generating realistic synthetic datasets. By capturing these relationships, BNs allow for the generation of synthetic data that not only mimics the distribution of the original data but also maintains the underlying causal structures, making it highly applicable for scenarios requiring interpretable and robust synthetic data [14].
    The process of generating synthetic data using Bayesian Networks involves the following steps [80]:
    1. (a)
      Structure Learning: The first step is to determine the structure of the Bayesian Network, which involves identifying the dependencies between variables. This can be done using expert knowledge or algorithms that learn the structure from data, such as hill climbing or score-based methods [81]. Each node in the network corresponds to a variable, and directed edges between nodes represent conditional dependencies.
    2. (b)
      Parameter Learning: After defining the structure, the next step is to estimate the conditional probability distributions (CPDs) for each variable in the network, given its parent node. This joint probability information is learned from the original dataset.
    3. (c)
      Sampling: After learning the structure and parameters (CPDs), synthetic data is generated by sampling from the Bayesian network. First, the root node is sampled from its marginal distribution, and then the child nodes are sampled based on their conditional probabilities. This process continues throughout the network until all variables are sampled.
    4. (d)
      Data Generation: The sampling process is repeated for the required number of synthetic data points. Each sample is a synthesized record for the dataset. This process ensures that the generated data depicts the dependencies learned from the original dataset.
  2. 2.
    Tabular Variational Autoencoder (TVAE) is a generative model designed to handle the complexities of tabular data, often including a mix of continuous and categorical variables. TVAE has the base idea of autoencoders that consists of two parts: an encoder network that learns to compress high-dimensional data into a low-dimensional, latent spacial representation (the code) and a decoder network that decompresses the compressed representation into the original domain as depicted in Figure 3 [82]. Thus, it extends the traditional Variational Autoencoder (VAE) framework by adapting it to the specific needs of tabular data, ensuring that the generated synthetic data accurately reflects the structure and distribution of the original dataset [83].
    Mathematical process of TVAE for Synthetic data generation:
    1. (a)
      Encoder-Decoder Architecture: TVAE uses an encoder-decoder structure where the encoder $q_{\phi}(z|x)$ maps the input data $x$ to a latent space $z$, and the decoder $p_{\theta}(x|z)$ reconstructs the data from this latent representation. The encoder outputs the parameters of a Gaussian distribution in the latent space:
      $q_{\phi}(z|x)=\mathcal{N}(z;\mu_{\phi}(x),\sigma_{\phi}(x)^{2}I)$ (1)
      Where $\mathcal{N}(z;\mu_{\phi}(x),\sigma_{\phi}(x)^{2}I)$ represents a multivariate normal (Gaussian) distribution where the latent variable $z$ is assumed to follow a Gaussian distribution with mean $\mu_{\phi}(x)$ and covariance matrix $\sigma_{\phi}(x)^{2}I$. The two parameters, $\mu_{\phi}(x)$ and $\sigma_{\phi}(x)$, are functions of the input $x$ and are learned by the encoder network. More precisely, $\mu_{\phi}(x)$ is the mean vector of the Gaussian distribution, parameterized by the input $x$. This vector represents the expected value of the latent variable $z$ for a given input $x$, determining the center of the distribution, and $\sigma_{\phi}(x)^{2}$ is the variance of the latent variable $z$, also parameterized by the input $x$. This variance determines how much the latent variable $z$ can deviate from its mean $\mu_{\phi}(x)$, controlling the spread of the distribution. Specifically, $\sigma_{\phi}(x)^{2}$ is the element-wise square of the standard deviation $\sigma_{\phi}(x)$. Plus, $I$ is the identity matrix used to represent that the latent variables are assumed to be independent, meaning there is no covariance between different latent dimensions. The use of $I$ ensures that the covariance matrix is diagonal, implying that each latent variable has its independent variance, controlled by $\sigma_{\phi}(x)^{2}$.
    2. (b)
      Variational Objective: The objective of training TVAE is to maximize the Evidence Lower Bound (ELBO):
      $\begin{split}\mathcal{L}(\theta,\phi;x)=&\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-\\ &D_{KL}(q_{\phi}(z|x)||p(z))\end{split}$ (2)
      Here, $\mathcal{L}(\theta,\phi;x)$ is the objective function (ELBO) to maximize during the training of TVAE. It measures how well the model reconstructs data and regularizes the latent space, and $\log p_{\theta}(x|z)$ represents the reconstruction loss, ensuring that the generated data $\hat{x}$ resembles the input data $x$. $D_{KL}$ is the Kullback-Leibler divergence that regularizes the latent space to follow a prior distribution $p(z)$, typically a standard normal distribution $\mathcal{N}(0,I)$. Plus, $\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]$ is the expected log-likelihood term, representing the reconstruction loss. It ensures that the generated data $\hat{x}$ resembles the input data $x$.
    3. (c)
      Latent Variable Modeling: The latent space $z$ is sampled from the Gaussian distribution parameterized by $\mu_{\phi}(x)$ and $\sigma_{\phi}(x)$. This latent space effectively captures the dependencies and distribution of the original data.
    4. (d)
      Data Generation: To generate synthetic data, samples are drawn from the latent space $z$, and these samples are passed through the decoder $p_{\theta}(x|z)$ to produce synthetic data points $\hat{x}$. The decoder ensures that the synthetic data preserves the statistical characteristics of the original tabular dataset.
    5. (e)
      Categorical and Continuous Data Handling: TVAE incorporates specific mechanisms to handle the peculiarities of tabular data. Standard Gaussian distributions are used for continuous variables, while TVAE often employs a Gumbel-Softmax distribution for categorical variables to allow gradient-based optimization.
    Figure 4: Diffusion Model Flow
  3. 3.
    Diffusion Models, particularly Denoising Diffusion Probabilistic Models (DDPM), have emerged as a powerful class of generative models [84, 85]. Originally developed for continuous data types such as images, these models have been adapted to handle the complexities of tabular data, which often includes a mix of categorical and numerical variables. The application of diffusion models to tabular data is exemplified by the Tabular Diffusion Model (TabDDPM), which has shown state-of-the-art performance in generating high-quality synthetic tabular data [86].
    Mathematical process of diffusion models for tabular data generation:
    1. (a)
      Forward and Reverse Diffusion Processes: The diffusion model begins with a forward process that gradually adds Gaussian noise to the data. Given a data point $x_{0}$, the forward process generates a sequence of increasingly noisy versions $x_{1},x_{2},\dots,x_{T}$, where each step adds a small amount of noise (Figure 4):
      $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (3)
      Here, $\beta_{t}$ is a variance schedule that controls the amount of noise added at each step $t$. The term $\sqrt{1-\beta_{t}}$ represents the decay factor, which ensures that the noise doesn’t overwhelm the signal, and $I$ is the identity matrix indicating that the noise is isotropic.
      The reverse process aims to generate synthetic data by denoising a sample starting from pure noise. The reverse distribution $p_{\theta}(x_{t-1}|x_{t})$ is parameterized by a neural network that predicts the mean and variance at each step:
      $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{\theta}^{2}I)$ (4)
      Here, $\mu_{\theta}(x_{t},t)$ is the predicted mean and $\sigma_{\theta}^{2}$ is the predicted variance at time step $t$, both parameterized by the neural network $\theta$. The network parameters $\theta$ are optimized to minimize the Kullback-Leibler divergence between the forward and reverse processes, effectively learning to denoise the data step by step.
    2. (b)
      Loss Function: The training objective for diffusion models in tabular data is typically the variational lower bound (VLB), which can be simplified to the mean-squared error (MSE) between the predicted noise and the actual noise:
      $\mathcal{L}(\theta)=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right]$ (5)
      Here, $\epsilon_{\theta}(x_{t},t)$ is the predicted noise by the model at time step $t$, where $x_{t}$ represents the noisy data and $\epsilon_{\theta}$ is parameterized by the neural network $\theta$. The term $\epsilon$ represents the noise added to the data during the forward process. The model minimizes the difference between $\epsilon$ and $\epsilon_{\theta}(x_{t},t)$, effectively learning to reverse the noise and reconstruct the data.
      The expectation $\mathbb{E}_{t,x_{0},\epsilon}$ denotes the averaging over all time steps $t$, initial data $x_{0}$, and noise samples $\epsilon$ applied during the forward process.
    3. (c)
      Handling Categorical and Numerical Features: Diffusion models for tabular data, such as TabDDPM, extend the basic framework to handle mixed data types. Categorical variables are typically processed using multinomial diffusion, where the data is represented as one-hot encoded vectors, and the noise is added in a way that respects the unconditional nature of the data:
      $q(x_{t}|x_{t-1})=\text{Cat}(x_{t};(1-\beta_{t})x_{t-1}+\beta_{t}/K)$ (6)
      Here, $\text{Cat}(x_{t};\cdot)$ represents the categorical distribution, which ensures that the synthetic data adheres to the unconditional nature of the variable. The term $\beta_{t}$ is the variance schedule controlling the amount of noise added, while $K$ represents the number of categories. $(1-\beta_{t})x_{t-1}+\beta_{t}/K$ maintains the correct distribution of categories by adjusting the probability mass of each category at every time step $t$.
      Numerical variables are handled using standard Gaussian diffusion processes. The model learns to denoise both categorical and numerical features in a unified framework, ensuring that the generated synthetic data maintains the relationships and distributions of the original data.
    4. (d)
      Synthetic Data Generation: After training, the model can generate synthetic tabular data by sampling from the learned distribution. The reverse process generates a sequence of less noisy samples from a sample of pure noise until the final synthetic data point is produced. This approach allows the model to create realistic tabular data that captures individual feature distributions and the dependencies between features.
    Figure 5: Original GAN structure by Goodfellow [9]
  4. 4.
    Generative Adversarial Networks (GAN)
Conditional Tabular GAN (CTGAN)
CTGAN is a generative model designed explicitly for tabular data, which is often characterized by the presence of mixed data types, including both categorical and continuous variables [12]. CTGAN extends the traditional GAN framework (Figure 5) by incorporating techniques that are well-suited for the unique challenges of tabular data generation, such as handling imbalanced datasets and preserving the relationships between variables [9, 3].
Mathematical process of CTGAN for synthetic data generation:
  1. 1.
    Generator and Discriminator Networks: CTGAN uses a generator network $G(z,c)$ that takes as input a noise vector $z$ sampled from a standard normal distribution and a conditional vector $c$ that encodes the categorical features. The generator produces synthetic data samples that mimic the real data distribution. The discriminator network $D(x,c)$ is then tasked with distinguishing between real data samples $x$ and synthetic samples generated by $G(z,c)$ while also considering the conditional vector $c$.
  2. 2.
    Mode-Specific Normalization for Continuous Data: To handle the diverse distribution of continuous features in tabular data, CTGAN introduces a mode-specific normalization technique. This approach transforms each continuous variable by centering it around a randomly sampled value from its distribution and normalizing it based on the empirical standard deviation of the data within that mode. This process ensures that the generator can more effectively model continuous data distributions:
    $x_{normalized}=\frac{x-x_{sample}}{\sigma(x_{mode})}$ (7)
    where $x$ is the original continuous variable, $x_{sample}$ is a value randomly sampled from the distribution of $x$, and $\sigma(x_{mode})$ is the empirical standard deviation of $x$ within its mode.
  3. 3.
    Conditional Vector Construction: CTGAN utilizes a conditional vector $c$ to guide the generation process, particularly for categorical variables. This vector is one-hot encoded, allowing the generator to focus on producing samples that match the specific conditions defined by $c$. The conditional vector is also integrated into the discriminator’s training process, helping it better differentiate between natural and synthetic data based on the specified conditions.
  4. 4.
    Training Objective: The training process of CTGAN follows the standard GAN framework, where the generator and discriminator are trained in a min-max game. The generator aims to minimize the following loss function to fool the discriminator, while the discriminator tries to maximize its ability to distinguish real from synthetic data:
    $\begin{split}\min_{G}\max_{D}&\mathbb{E}_{x\sim p_{data}}[\log D(x,c)]+\\ &\mathbb{E}_{z\sim p_{z}}[\log(1-D(G(z,c),c))]\end{split}$ (8)
    Where $G$ represents the generator, $D$ represents the discriminator, $x$ is a data sample drawn from the real data distribution $p_{data}$, $z$ is a random noise vector drawn from the prior distribution $p_{z}$, $c$ denotes any additional conditional information, $G(z,c)$ does the generator generate the synthetic data, and $D(x,c)$ and $D(G(z,c),c)$ are the discriminator’s outputs for real and synthetic data, respectively.
  5. 5.
    Data Sampling and Imbalance Handling: CTGAN employs a training-by-sampling technique to address class imbalance in tabular data. The data points are sampled during training to ensure underrepresented classes are adequately represented in the training batches. This approach mitigates the impact of imbalanced datasets on the quality of the generated data.
Copula GAN
CopulaGAN is a generative model that combines the strengths of copula-based statistical models with the flexibility of GANs to generate synthetic tabular data [87]. Copulas are mathematical tools that describe the dependence structure between random variables, enabling the modeling of complex dependencies separately from the marginal distributions [88]. By integrating copulas into the GAN framework, Copula GAN effectively captures the intricate dependencies in tabular data, making it particularly suitable for scenarios where preserving the relationships between variables is crucial [89].
The mathematical process of Copula-GAN for synthetic data generation:
  1. 1.
    Marginal Distribution Estimation: The first step in Copula GAN involves estimating the marginal distributions of each variable in the tabular dataset. This can be achieved using kernel density estimation for continuous variables or empirical distributions for categorical variables. The goal is to transform the original data into a uniform distribution on the interval $[0,1]$ using the cumulative distribution function (CDF) of each variable:
    $u_{i}=F_{i}(x_{i})\quad\text{for }i=1,\dots,d$ (9)
    where $u_{i}$ is the transformed variable, and $F_{i}(x_{i})$ is the CDF of the $i$-th variable.
  2. 2.
    Copula Function for Dependency Modeling: Copula GAN models the dependencies between the variables using a copula function once the data is transformed into the copula space. The copula function $C$ captures the joint distribution of the transformed variables:
    $C(u_{1},\dots,u_{d})=\mathbb{P}(U_{1}\leq u_{1},\dots,U_{d}\leq u_{d})$ (10)
    where $C$ denotes the copula function, $u_{1},\dots,u_{d}$ are the values of the copula-transformed variables, and $U_{1},\dots,U_{d}$ are the random variables corresponding to these transformed values. The expression $\mathbb{P}(U_{1}\leq u_{1},\dots,U_{d}\leq u_{d})$ represents the joint probability that $U_{1}$ is less than or equal to $u_{1}$, $U_{2}$ is less than or equal to $u_{2}$, and so on up to $U_{d}$ and $u_{d}$.
    The copula function separates the dependency structure from the marginal distributions, which is advantageous for generating synthetic data that preserves the original data’s relationships.
  3. 3.
    Generator and Discriminator Networks: In Copula GAN, the generator $G(z)$ takes a noise vector $z$ and generates samples in the copula space, aiming to mimic the dependency structure captured by the copula function. The discriminator $D(u)$ evaluates whether a given sample from the copula space is real or generated, guiding the generator to produce more realistic samples:
    $\begin{split}\min_{G}\max_{D}&\mathbb{E}_{u\sim p_{data}}[\log D(u)]+\\ &\mathbb{E}_{z\sim p_{z}}[\log(1-D(G(z)))]\end{split}$ (11)
  4. 4.
    Inverse Transformation and Synthetic Data Generation: After training, the generator produces samples in the copula space. These samples are then transformed back to the original data space using the inverse of the marginal CDFs:
    $x_{i}=F_{i}^{-1}(u_{i})\quad\text{for }i=1,\dots,d$ (12)
    Where $x_{i}$ represents the $i$-th component of the synthetic data in the original data space, $u_{i}$ is the $i$-th element in the sample in the copula space, and $F_{i}^{-1}$ denotes the inverse of the marginal CDF of the $i$-th variable. This inverse transformation ensures that the generated data points follow the original marginal distributions while maintaining the dependencies modeled by the copula.
Naive Bayes and Logistic Regression GAN (GAN- BLR++
GANBLR++ is an advanced generative model designed to address the limitations of traditional GANs in generating synthetic tabular data, mainly where complex dependencies and mixed data types (categorical and numerical) are involved. Unlike its predecessor, GANBLR, which primarily focused on modeling categorical data using restricted Bayesian Networks, GANBLR++ introduces significant enhancements that allow it to generate both categorical and numerical data with greater accuracy [14, 90]. This model integrates unrestricted Bayesian Networks and introduces a Dirichlet Mixture Model (DMM) for numerical data generation, enabling it to capture the full range of dependencies in real-world tabular datasets [91]. GANBLR++ is particularly useful in domains where the integrity of the generated data is critical, such as cybersecurity, which offers a robust solution that balances flexibility, accuracy, and computational efficiency.
The mathematical process of GANBLR++ for synthetic data generation:
  1. 1.
    Incorporation of Unrestricted Bayesian Networks: GANBLR++ leverages unrestricted Bayesian Networks to model complex dependencies between variables. The network structure is learned through a scoring function that optimizes the log-likelihood of the data while including a regularization term to avoid overfitting. This allows the model to represent more intricate relationships between variables than the restricted networks used in GANBLR.
    $\text{Score}(G:D_{\text{data}})=\text{LL}(G:D_{\text{data}})-\lambda\cdot|G|$ (13)
    where $\text{Score}(G:D_{data})$ is the score of the network structure $G$ given the data $D_{data}$, $\text{LL}(G:D_{data})$ denotes the log-likelihood of the data under the network $G$, $\lambda$ is the regularization parameter, and $|G|$ represents the complexity of the network $G$. The regularization term $\lambda\cdot|G|$ helps prevent overfitting by penalizing more complex network structures.
  2. 2.
    Dirichlet Mixture Model for Numerical Data: To handle numerical data effectively, GANBLR++ uses a Dirichlet Mixture Model (DMM). The DMM identifies modes within the numerical data distribution and samples from these modes during data generation, ensuring that the generated values accurately reflect the underlying distribution of the original data.
    $x_{i}\sim\text{N}(\mu_{c},\sigma_{c})$ (14)
    where $x_{i}$ is the $i$-th sample from the numerical data, $\text{N}(\mu_{c},\sigma_{c})$ denotes a Gaussian distribution with mean $\mu_{c}$ and standard deviation $\sigma_{c}$ for a particular mode $c$. The parameters $\mu_{c}$ and $\sigma_{c}$ represent the mean and standard deviation of the Gaussian distribution associated with the mode $c$.
  3. 3.
    Mode-Specific Sampling: GANBLR++ further refines the generation of numerical data through mode-specific sampling, which ensures that the generated data points not only match the distributional characteristics of the original data but also maintain the necessary granularity.
    $\begin{split}&x_{i}\sim\text{N}(\mu_{c},\sigma_{c})\quad\text{subject to }\\ &\alpha_{c}<\frac{x_{i}-\mu_{c}}{\sigma_{c}}\leq\alpha_{c}+1\end{split}$ (15)
    where $x_{i}$ is the $i$-th sample from the numerical data, $\text{N}(\mu_{c},\sigma_{c})$ denotes a Gaussian distribution with mean $\mu_{c}$ and standard deviation $\sigma_{c}$ for a particular mode $c$, and $\alpha_{c}$ is a mode-specific threshold parameter. The constraint $\alpha_{c}<\frac{x_{i}-\mu_{c}}{\sigma_{c}}\leq\alpha_{c}+1$ ensures that $x_{i}$ falls within a specific range relative to the mode $c$, maintaining the granularity of the distribution.
Cascaded Tabular GAN (CasTGAN)
CasTGAN is a novel framework designed for generating realistic tabular data [92]. Traditional GAN models often struggle with capturing the complex dependencies between features in tabular datasets, leading to synthetic data that fails to preserve the validity and interdependencies present in the original data [93]. CasTGAN addresses this limitation by employing a cascaded architecture where each generator in the sequence is responsible for synthesizing a single feature of the dataset. This architecture allows the model to better capture the intricate relationships between features, ensuring that the generated data is not only statistically similar to the real data but also semantically valid. The primary motivation behind CasTGAN is to enhance the generation of high-quality synthetic tabular data, which is crucial in scenarios where data privacy is a concern and real data cannot be freely shared or used.
The mathematical framework of CasTGAN can be broken down into the following steps:
  1. 1.
    Cascaded Generators: The model is composed of multiple generators $G_{1},G_{2},\dots,G_{M}$, each responsible for generating one feature of the dataset. Each generator $G_{i}$ receives inputs from a noise vector $z$ and the outputs of the previous generator $G_{i-1}$.
  2. 2.
    Auxiliary Learners: For each generator $G_{i}$, there is an auxiliary learner $AL_{i}$ that computes a loss specific to the feature generated by $G_{i}$. The auxiliary loss guides the generator toward producing more accurate and valid feature values.
  3. 3.
    Discriminator: A single discriminator $D$ evaluates the final synthetic dataset, distinguishing between real and generated data. The discriminator is trained on the output of the final generator $G_{M}$, and its loss is propagated back through the cascade to improve the performance of all generators.
  4. 4.
    Training Objective: The training of CASTGAN involves minimizing the Wasserstein distance with a gradient penalty between the real and generated data distributions. The value function for the discriminator and the generators is given by:
    $\begin{split}&\min_{G_{\rightarrow}}\max_{D}V(D)=\mathbb{E}_{x\sim p_{data}}[D(x)]\\ &-\mathbb{E}_{z\sim p_{z}}[D(G_{\rightarrow}(z))]\\ &-\lambda_{GP}\mathbb{E}_{\hat{x}\sim p_{\hat{x}}}\left[(\|\nabla_{\hat{x}}D(\hat{x})\|_{2}-1)^{2}\right]\end{split}$ (16)
    where $G_{\rightarrow}$ represents the cascade of generators, $D$ is the discriminator, $x$ is a sample from the real data distribution $p_{data}$, $z$ is a sample from the prior distribution $p_{z}$, $G_{\rightarrow}(z)$ is the generated sample from the cascade of generators, $\hat{x}$ is a sample from the data distribution used for the gradient penalty $p_{\hat{x}}$, $\nabla_{\hat{x}}D(\hat{x})$ is the gradient of the discriminator concerning $\hat{x}$, $\|\nabla_{\hat{x}}D(\hat{x})\|_{2}$ denotes the L2 norm of this gradient, and $\lambda_{GP}$ is the gradient penalty coefficient. The term $(\|\nabla_{\hat{x}}D(\hat{x})\|_{2}-1)^{2}$ enforces the Lipschitz constraint on the discriminator.
  5. 5.
    Auxiliary Loss Integration: Each generator $G_{i}$ minimizes a combination of the discriminator loss and its corresponding auxiliary loss, scaled by a coefficient $\lambda_{AL_{i}}$.
  6. 6.
    Final Output: The final synthetic dataset $\hat{X}$ is generated by passing a noise vector $z$ through the entire cascade of generators.
IV Mutual Informatiaon
Mutual Information (MI) measures the mutual dependence between two random variables (linear or non-linear), initially introduced by Shannon in his foundational work on information theory [94]. It quantifies the amount of information gained about one random variable through the observation of another [95]. Unlike the correlation coefficient, limited to real-valued random variables and linear relationships, MI is more general, capturing both linear and non-linear dependencies [96]. The concept of MI is closely related to entropy, a fundamental notion in information theory that measures the expected amount of information in a random variable [97]. MI can be understood as the expected value of the pointwise mutual information (PMI), reflecting how much the joint distribution of two variables deviates from the product of their marginal distributions [98].
IV-A Comparison with Other Methods
The rationale for selecting MI over other methods lies in its comprehensive mathematical framework. Initially, this study considered two general approaches for choosing the best features based on their dependence on the target feature and their relationships. Correlation was the first and most obvious choice, mathematically defined as the Pearson correlation coefficient (equ: LABEL:EQ:_CORR). However, correlation only accounts for linear relationships [99]. In real-world scenarios, non-linear relationships often need to be identified and quantified to select the optimal set of features.
$\text{Corr}(X,Y)=\frac{\text{Cov}(X,Y)}{\sigma_{X}\sigma_{Y}}$ (17)
where $\text{Cov}(X,Y)$ is the covariance between variables $X$ and $Y$, and $\sigma_{X}$ and $\sigma_{Y}$ are the standard deviations of $X$ and $Y$, respectively. While correlation is effective for capturing linear relationships, it does not account for the non-linear relationships that often exist in real-world data, limiting its effectiveness for feature selection.
The second option was a tree-based AI technique, known for its robustness and ability to identify essential features rigorously [100]. However, tree-based methods suffer from the ”black box” problem, offering little interpretability regarding why certain features are selected.
While approaches like Information Gain (IG) have been used for feature selection, as demonstrated by one study [101], Information Gain only measures the relationship between individual features and the target variable, overlooking interdependencies between features [102]. Mathematically, it measures the reduction in entropy when a feature $X$ is used to predict the target variable $Y$.
$IG(Y,X)=H(Y)-H(Y|X)$ (18)
where $H(Y)$ is the entropy of the target variable $Y$, and $H(Y|X)$ is the conditional entropy of $Y$ given $X$. Although Information Gain quantifies the relationship between individual features and the target variable, it ignores interdependencies among the features.
In contrast, MI captures the relationship between features and the target and the dependencies among the features themselves, providing a more holistic evaluation [103]. MI is defined as:
$MI(X,Y)=H(X)+H(Y)-H(X,Y)$ (19)
where $H(X)$ and $H(Y)$ are the entropies of variables $X$ and $Y$, and $H(X,Y)$ is their joint entropy. By capturing non-linear dependencies and interactions among features, MI offers a more comprehensive framework for feature selection.
Compared to other AI-based feature selection methods, such as recursive feature elimination or embedded methods based on decision trees [104], MI has the upper hand by being model-agnostic. Many AI-based techniques are tied to specific algorithms or models, which may introduce biases or limit generalizability across different datasets or classifiers. MI, however, evaluates feature importance based on statistical dependencies independent of any specific model, making it a versatile and unbiased approach [105]. This allows MI to provide a more robust selection process, especially in scenarios where feature relationships are complex and nonlinear, a common occurrence in cybersecurity network data.
IV-B Application of Mutual Information in this Study
In this study, MI is crucial for measuring the dataset’s dependency between different features (variables). It helps understand the strength of relationships between variables, which is essential for modeling and analysis. For example, mutual information was used to quantify the relation between each feature $x_{i}$ and the target $y$ and between each feature to identify interdependencies within the features. This step was crucial to filter down the essential features that explain the target variables and to cluster the features exhibiting similar behavior.
The mutual information score of every feature was determined concerning the target, and the top 25% of the features were selected based on the mutual information weights (Table III & Table IV). This way, the computation was less intensive, the information was high, and the results were more interpretable.
V Evaluation Metrics
The evaluation metrics utilized in this study align with those used in previous studies, as outlined in section II of this paper. The following section details how these metrics were employed in this study and how each metric was measured to assess the experimental results.
  1. 1.
    Statistical Similarity (Fidelity) refers to the necessity for synthetic data to resemble the underlying statistical properties of real data closely [106]. This metric’s core mathematical definition is that each variable’s probability distribution in the synthetic dataset should closely match that of the corresponding variable in the real dataset [44]. However, beyond replicating the individual behavior of each variable, it is equally important to examine the interdependencies and relationships between variables. To evaluate these relationships, we compare the correlations among variables in both datasets [107].
    Another crucial aspect is the preservation of the data structure. For instance, if a variable is binary in the real data, it must remain binary in the synthetic data to ensure logical consistency. Similarly, suppose a variable is defined to have values greater than or equal to zero (such as packet size in network traffic data). In that case, the corresponding variable in the synthetic data should also respect this constraint. Failing to maintain these structural properties can lead to synthetic data that, while statistically similar on a superficial level, fails to capture the logical and contextual nuances necessary for accurate downstream analyses and applications [108]. In this study, the following metrics were used to measure the statistical similarity (fidelity), as shown in Tables V to X:
    1. (a)
      Data Boundaries (DB): This metric checks whether the synthetic data adheres to the original data’s logical minimum and maximum values. For example, a binary column should contain only binary values, ensuring no out-of-bound entries.
    2. (b)
      Correlation (Corr): This is assessed by examining the correlation heatmap of all variables’ absolute values of the correlation coefficients. It evaluates the strength and direction of relationships in both real and synthetic datasets.
    3. (c)
      Probability Distribution (PD): This compares the underlying probability distribution of each variable between the real and synthetic datasets, ensuring that their statistical properties are aligned.
  2. 2.
    Performance (utility/accuracy) is a widely used metric to evaluate the utility of synthetic data, particularly in the context of ML and DL models [109]. This metric compares how models trained on synthetic data perform relative to those trained on real data [110]. The most common approach is to contrast TRTR (train and test on real data) with TSTR (train on synthetic data and test on real data). By assessing the model’s accuracy, F1 score, precision, and recall, researchers can determine the efficacy of the synthetic data. Suppose the model’s performance metrics in the TSTR scenario are comparable to or exceed those in the TRTR scenario. In that case, the synthetic data is considered a viable alternative to the real data [14].
    Moreover, a well-performing TSTR indicates that the synthetic data captures the essential patterns and relationships within the real data, making it suitable for model development, validation, and deployment in privacy-sensitive environments. However, it’s crucial to note that synthetic data’s utility is not solely dependent on performance metrics; it also relies on the data’s ability to generalize across different models and tasks, which should be considered in comprehensive evaluations [111].
    Thus, in this study, the accuracy of TRTR was compared with TSTR for measuring ML utility, as reported in Tables V to X.
  3. 3.
    Class Balance (CB) is a critical metric for evaluating the quality of synthetic tabular data, particularly in classification tasks. Class balance refers to the distribution of instances across different classes in the dataset[112]. Maintaining a balanced class distribution in real-world datasets is essential for training robust ML models, as imbalanced courses can lead to biased models that perform poorly on underrepresented classes [113].
    Ensuring that the class distribution closely mirrors the real data is essential when assessing synthetic data. This involves not only comparing the overall proportions of each class but also examining the stability of these proportions across different subsets of the data. Some methods and tools for generating synthetic data inherently include mechanisms to evaluate and maintain class balance, ensuring that the synthetic dataset accurately reflects the class distribution of the real dataset. Suppose a synthetic data generation method does not inherently address class balance. In that case, it may require additional post-processing to adjust class distributions and avoid introducing bias into the model training process [114].
    Thus, the CB results mentioned in Tables V to X represent how much difference exists between the NORMAL and ATTACK cases. If the classes in the synthetic data are the same, there is zero difference; otherwise, the difference is detailed in the results table later in this paper.
VI Experiment Setup
TABLE I: Summary of 26 Encoded selected Features from NSL-KDD Dataset
S.No Feature Description
1 src_bytes Number of data bytes from source to destination
2 dst_bytes Number of data bytes from destination to source
3 same_srv_rate Percentage of connections to the same service
4 diff_srv_rate Percentage of connections to different services
5 flag_SF Connection status with a normal connection (SF: ”Normal”)
6 dst_host_srv_count Number of connections to the same service as the current connection in the past 100 connections
7 dst_host_same_srv_rate Percentage of connections to the same service for a destination host
8 logged_in 1 if successfully logged in; 0 otherwise
9 dst_host_serror_rate Percentage of connections that have “SYN” errors
10 dst_host_diff_srv_rate Percentage of connections to different services for a destination host
11 dst_host_srv_serror_rate Percentage of connections that have “SYN” errors for a destination host
12 serror_rate Percentage of connections that have “SYN” errors
13 srv_serror_rate Percentage of connections that have “SYN” errors for the same service
14 flag_S0 Connection status where no data packets were exchanged (S0: ”No Data Exchange”)
15 count Number of connections to the same host as the current connection in the past 2 seconds
16 service_http HTTP service (1 if used, 0 otherwise)
17 dst_host_srv_diff_host_rate Percentage of connections to different hosts on the same service
18 level Threat level of the connection
19 dst_host_count Number of connections to the same destination host in the past 100 connections
20 dst_host_same_src_port_rate Percentage of connections with the same source port to the destination host
21 service_private Private network service (1 if used, 0 otherwise)
22 srv_diff_host_rate Percentage of connections to different hosts for the same service
23 srv_count Number of connections to the same service as the current connection in the past 2 seconds
24 dst_host_srv_rerror_rate Percentage of connections that have “REJ” errors for a destination host
25 service_domain_u Domain name service (DNS) (1 if used, 0 otherwise)
26 target Class label indicating if the connection is normal or an attack
TABLE II: Summary of 21 selected Features from CICIDS-17 Dataset
S.No Feature Description
1 Average Packet Size Average size of the packets in the flow
2 Packet Length Std Standard deviation of the packet lengths in the flow
3 Packet Length Variance Variance of the packet lengths in the flow
4 Packet Length Mean Mean length of the packets in the flow
5 Total Length of Bwd Packets Total number of bytes in backward (Bwd) packets
6 Subflow Bwd Bytes Number of bytes in the backward sub-flow
7 Destination Port Port number of the destination host
8 Avg Bwd Segment Size Average size of backward segments
9 Bwd Packet Length Mean Mean length of backward packets
10 Init_Win_bytes_forward Initial window size in bytes for forward direction
11 Subflow Fwd Bytes Number of bytes in the forward sub-flow
12 Total Length of Fwd Packets Total number of bytes in forward (Fwd) packets
13 Max Packet Length Maximum length of a packet in the flow
14 Bwd Packet Length Max Maximum length of a backward packet
15 Init_Win_bytes_backward Initial window size in bytes for backward direction
16 Fwd Packet Length Max Maximum length of a forward packet
17 Fwd Packet Length Mean Mean length of forward packets
18 Avg Fwd Segment Size Average size of forward segments
19 Flow IAT Max Maximum time interval between packets in the flow
20 Flow Bytes/s Rate of flow in bytes per second
21 target Class label indicating if the flow is benign or malicious
TABLE III: Mutual Information Score of NSL-KDD Features
Feature Mutual Information Score
src_bytes 0.566864
dst_bytes 0.439281
same_srv_rate 0.369288
diff_srv_rate 0.361895
flag_SF 0.341828
dst_host_srv_count 0.335993
dst_host_same_srv_rate 0.309832
logged_in 0.292075
dst_host_serror_rate 0.286589
dst_host_diff_srv_rate 0.283874
dst_host_srv_serror_rate 0.281332
serror_rate 0.278666
srv_serror_rate 0.269186
flag_S0 0.263399
count 0.262733
service_http 0.191343
dst_host_srv_diff_host_rate 0.189822
level 0.153819
dst_host_count 0.144479
dst_host_same_src_port_rate 0.131316
service_private 0.118493
srv_diff_host_rate 0.099441
srv_count 0.062476
dst_host_srv_rerror_rate 0.062244
service_domain_u 0.048430
TABLE IV: Mutual Information Score of CICIDS-17 Features
Feature Mutual Information Score
Average Packet Size 0.347112
Packet Length Std 0.342188
Packet Length Variance 0.342019
Packet Length Mean 0.319635
Total Length of Bwd Packets 0.296955
Subflow Bwd Bytes 0.296882
Destination Port 0.291245
Avg Bwd Segment Size 0.287676
Bwd Packet Length Mean 0.287483
Init_Win_bytes_forward 0.287174
Subflow Fwd Bytes 0.284528
Total Length of Fwd Packets 0.284264
Max Packet Length 0.264060
Bwd Packet Length Max 0.263210
Init_Win_bytes_backward 0.250188
Fwd Packet Length Max 0.246537
Fwd Packet Length Mean 0.213223
Avg Fwd Segment Size 0.213065
Flow IAT Max 0.212817
Flow Bytes/s 0.209784
VI-A Hardware Setup
The experiments were performed on a workstation with a 13th Gen Intel® Core™ i9-13900 processor. This processor operates at 2000 MHz and features 24 cores with 32 logical processors. The system is equipped with 32 GB of RAM. A dedicated GPU, NVIDIA GeForce RTX 4090, was used for these experiments.
VI-B Software Setup
The software environment for the experiments was based on Microsoft Windows 11 Pro. The primary programming language used was Python 3.12.4, which provides a stable and versatile platform for implementing various data science and ML techniques. Essential libraries and frameworks include NumPy for numerical operations, Pandas for data manipulation, Seaborn and matplotlib for data visualization, and sci-kit-learn for ML utilities. The experiments also leveraged PyTorch for deep learning tasks. For handling class imbalance, the imbalanced-learn library (learn) was employed. The libraries are crucial for the GAN models, including SDV (Synthetic Data Vault) for tabular data generation, as they support models like CTGAN, TVAE, and CopulaGAN. Due to the lack of pre-existing packages, their official GitHub repositories were also utilized for specialized GAN models such as GANBLR and CASTGAN. PGMPy and SciPy were also used for probabilistic graphical models and scientific computations.
VI-C Datasets & Pre-processing
  1. 1.
    NSL-KDD dataset, a refined version of the original KDD Cup 1999 dataset, is widely used in cybersecurity to evaluate IDS [115]. It addresses some critical issues present in its predecessor, such as redundant records and imbalanced distribution, thereby providing a more accurate and reliable benchmark for performance evaluation. For this study, the training portion of the NSL-KDD dataset, originally consisting of 125,973 instances and 42 columns (41 features and 1 (binary) target), was used.
    The preprocessing of the NSL-KDD dataset involved several essential steps to prepare it for synthetic data generation and evaluation. Initially, categorical columns were encoded to transform non-numeric data into a suitable format for ML algorithms. This encoding process ensured that all categorical variables were converted into numerical values, facilitating the subsequent analysis. Following the encoding, feature extraction was performed to identify the most significant features using Mutual Information. This process selected the features’ top 25% (Q1) based on the mutual information weights. This step resulted in a reduced feature set of 26 encoded features (Table I), which were the most important (based on the information weights; Table III) for accurately representing the data. This refined feature set was then used in the data generation phase, ensuring that the synthetic data closely mirrored the critical characteristics of the original dataset.
  2. 2.
    CICIDS-17 dataset, part of the Canadian Institute for Cybersecurity’s intrusion detection dataset collection, is widely employed for evaluating cybersecurity systems [18]. It provides a comprehensive network traffic data set with diverse attack types and normal traffic, offering a robust benchmark for performance assessment. CICIDS-17 has one week of captured network traffic data in eight comma-separated files (CSV). These files are generated based on the attack types. There are 2,830,743 instances of all eight files and 79 columns (78 numerical features and one binary target).
    Preprocessing of the CICIDS-17 dataset involved several vital steps to ready it for synthetic data generation and evaluation. First, all the files were individually tested for null values and duplicated columns. In this process, it was found that there were some infinity values in the data. Thus, all the null values, infinity, and duplicated column ”Fwd Header Length” were removed. Secondly, all the cleaned data files were concatenated into a single CSV file for further feature selection. Lastly, mutual information was performed to select the variables’ top 25% (Q1) based on the mutual information weights. This step resulted in a reduced feature set of 21 features (Table II), which were determined to be the most important (based on the information weights; IV) for accurately representing the data. This reduced feature set was then employed in the synthetic data generation phase, ensuring that the synthetic data accurately reflected the essential characteristics of the original dataset.
VII Experimental Results and Discussion
TABLE V: Comparison of Synthetic Data by Standard Methods on NSL-KDD Dataset
Method Description Pros Cons Performance Summary
SS CB Accuracy
DS Corr PD
ROS [116] Randomly duplicates minority class samples Simple, low cost Overfitting risk, no new samples Yes Yes 0% diff 0% diff 0.9998
SMOTE [63] Generates synthetic samples by interpolation Reduces overfitting, better minority representation May create ambiguous samples, high computation No Yes 0% diff 0% diff 0.9999
Adaptive Synthetic Sampling [67] Adjusts sampling based on data distribution Flexible, effective for imbalance Complex tuning, high computation No Yes 3.8% diff 0.14% diff 0.9998
Cluster based oversampling [117] Oversamples within identified clusters Captures local structures Sensitive to clustering algorithm No Yes 0% diff 0% diff 0.9995
Gaussian Mixture Model [75] Uses Gaussian distributions for sampling Captures complex distributions Assumes data distribution No Yes 99.1% diff 99.9% diff 0.5314
SS” = Statistical Similarity, ”DS” = Data Structure, ”Corr” = Correlation, ”PD” = Probability Distribution, ”CB” = Class Balance, ”Accuracy” = Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data = 0.9995, ”TSTR” = Train on Synthetic and Test on Real Data
TABLE VI: Comparison of Synthetic Data by Standard Methods on CICIDS-17 Dataset
Method Statistical Similarity Class Balance Accuracy (0.9978)
Data Structure Correlation Probability Distribution
ROS Yes Yes 0% diff 0% diff 0.9991
SMOTE Yes Yes 0% diff 0% diff 0.9990
Adaptive Synthetic Sampling Yes Yes 0% diff 0.2% diff 0.9986
Cluster-based oversampling No Yes 0% diff 0% diff 0.9930
Gaussian Mixture Model No No 61.9% diff 46.1% diff 0.4509
Accuracy” = Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
TABLE VII: Comparison of Synthetic Data by AI Methods on NSL-KDD Dataset
Method Description Pros Cons Performance Summary
SS CB Accuracy
DS Corr PD
Bayesian Neural Networks (BN) [118] Uses Bayesian inference for uncertainty estimation Quantifies uncertainty, probabilistic dependencies Requires prior knowledge, complex tuning No No 0% diff 17.2% diff 0.9587
Variational Tabular Autoencoders (TVAEs)[119] Neural networks for unsupervised feature learning Learns representations, reduces dimensionality KL divergence issue, risk of overfitting Yes Yes 23.1% diff 26.7% diff 0.9807
Diffusion Models [86] Uses diffusion for data generation Captures temporal dependencies Needs sequential data, high computational cost No No 98.3% diff 34.2% diff 0.51
Conditional Tabular Generative Adversarial Networks (CTGAN) [12] Adversarial training between generator and discriminator High-quality data, models complex distributions Training instability, mode collapse Yes Yes 11.5% diff 5.7% diff 0.9640
SS = Statistical Similarity, DS = Data Structure, Corr = Correlation, PD = Probability Distribution, CB = Class Balance, ”Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data = 0.9995, ”TSTR” = Train on Synthetic and Test on Real Data
TABLE VIII: Comparison of Synthetic Data by AI methods on CICIDS-17 Dataset
Method Statistical Similarity Class Balance Accuracy (0.9978)
Data Structure Correlation Probability Distribution
Bayesian Neural Networks (BN)
Tabular Variational Autoencoder (TVAE) Yes No 1.4% diff 69.4% diff 0.9718
Conditional Tabular GAN (CTGAN) Yes Yes 0.8% diff 5.2% diff 0.9603
Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
TABLE IX: Comparison of Synthetic Data by prominent GANs on NSL-KDD Dataset
Method Statistical Similarity Class Balance Accuracy (0.9995)
Data Structure Correlation Probability Distribution
CTGAN Yes Yes 7.7% diff 6.7% diff 0.9667
GANBLR++ No No 99% diff 8.8% diff 0.5916
CopulaGAN Yes Yes 7.7% diff 5.4% diff 0.9761
CastGAN No Yes 27% diff 11.1% diff 0.9582
Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
TABLE X: Comparison of Synthetic Data by prominent GANs on CICIDS-17 Dataset
Method Statistical Similarity Class Balance Accuracy (0.9978)
Data Structure Correlation Probability Distribution
CTGAN Yes Yes 0% diff 5% diff 0.95381
GANBLR++ No No 47.6% diff 64.4% diff 0.5378
CopulaGAN Yes Yes 0% diff 5% diff 0.9755
CastGAN No No 4.8% diff 58.3% diff 0.9734
Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
Given the surge in interest in generating synthetic tabular data, the GAN framework has emerged as a seminal approach, attracting substantial research attention. Consequently, this paper delves deeply into evaluating current high-performing GAN models, leveraging recent advancements in this domain. A meticulous analysis of four leading models is presented in Section III, elucidating their methodologies and performance metrics.
VII-A Results
This study evaluated various methods for generating synthetic tabular data, covering both statistical and AI-based approaches.
VII-A1 Statistical Methods
Five widely recognized statistical methods were assessed, as outlined in the survey by [3] (detailed in Section III). The findings indicate the following:
  • ROS emerged as the most effective statistical method for preserving the data distribution and ensuring class balance. It maintained a high degree of statistical similarity to the original dataset.
  • SMOTE and Cluster Centroid outperformed ROS by maintaining statistical similarity and enhancing ML utility, achieving a $99\%$ accuracy while preserving class balance.
  • GMM showed strong performance in capturing feature correlations but underperformed in statistical similarity and ML utility, with an accuracy of only $53\%$.
  • The results were consistent across both the synthetic NSL-KDD and CICIDS-17 datasets (Table V and Table VI).
VII-A2 AI Methods
Four prominent AI and generative AI methodologies were evaluated (details in Section III). The results are summarized as follows:
  • Bayesian Networks struggled with maintaining correlation and structure in the NSL-KDD dataset, especially for binary features. Diffusion models had the lowest ML utility accuracy at $51\%$.
  • TVAE and Conditional Tabular CTGAN performed exceptionally well, preserving statistical properties and achieving high ML utility accuracy rates of $98\%$ and $96\%$, respectively (Table VII).
  • Consistent performance was observed with the CICIDS-17 dataset, where TVAE slightly outperformed CTGAN in ML utility.
VII-A3 Prominent GANs
Four leading GAN models were evaluated in depth (detailed in Section III). The results showed:
  • CTGAN and Copula-GAN outperformed other models, with ML utility accuracy scores of $97\%$ and $98\%$, respectively, on the NSL-KDD and CICIDS-17 datasets.
  • CastGAN showed strong performance in handling feature dependencies but struggled to preserve boundary conditions for specific binary columns.
  • GANBLR++, based on Bayes’ network principles, exhibited performance metrics closely aligned with those of individual BN models.
VII-B Discussion
The results from this study provide several key insights into the effectiveness of various synthetic data generation methods:
VII-B1 Statistical Methods
While effective at maintaining data distribution, ROS is limited in generating genuinely synthetic data, making it less suitable for cybersecurity applications. By definition, ROS does not generate novel data points; it is just random sampling. SMOTE and Cluster Centroid, by enhancing ML utility, present better alternatives for applications where both statistical fidelity and predictive accuracy are critical. Both SMOTE and cluster centroids create novel data points by interpolation and average representation, respectively [63, 120]. GMM’s performance highlights its strength in capturing feature correlations and creating novel data points using the fundamental definition of samples from fitted distributions. Still, its lower utility score suggests it may not be ideal for all applications [121, 122].
VII-B2 AI Methods
Bayesian Networks’ struggle with larger datasets, particularly with multi-class features, indicates a need for better handling such data structures in AI models. The potential reason for this struggle comes from the probabilistic graphical model. This graph tries to represent dependency between variables, and in the case of the CIC-IDS large dataset, it fails to create discretization because of low representation learning (memory) [123, 124]. Diffusion models require further refinement to improve their statistical similarity and utility because they are, in general, designed for high-dimensional data with spatial relationships (e.g., images), making it challenging to capture the non-spatial, relational structure of tabular data [85]. The success of TVAE and CTGAN emphasizes the growing importance of generative AI models in synthetic data generation. However, their lack of inherent class imbalance solutions suggests that additional techniques may be required to address this issue comprehensively.
TVAE and CTGAN excel in generating high-quality synthetic tabular data due to their specialized architectures and loss functions. TVAE uses a probabilistic latent space to model complex dependencies, handling continuous and categorical features with different probability distributions [125]. Its KL divergence-based loss regularizes the model, ensuring diverse yet realistic data generation [126]. CTGAN, on the other hand, uses a conditional generator to capture feature relationships and a specific loss function that optimizes for mode-specific normalization, preserving the distributions of both majority and minority categories [127]. The combination of these tailored architectures and well-designed loss functions enables both models to generate synthetic data that closely mirrors the statistical properties of the original dataset, making them highly effective for tabular data generation [12].
VII-B3 Prominent GANs
The superior performance of CTGAN and Copula-GAN reinforces their suitability for generating synthetic tabular data that retains statistical integrity and practical utility. CopulaGAN excels in generating synthetic tabular data because it uses copulas to model the dependencies between variables and their marginal distributions separately [128]. This allows it to capture complex relationships between features more accurately than standard GANs. By decoupling the modeling of individual feature distributions from their joint dependency structure, CopulaGAN ensures that the synthetic data maintains the statistical properties of the original dataset, making it highly effective for tabular data [129]. CastGAN’s innovative architecture of each generator and auxiliary learner for each variable shows promise, though its limitations with binary features suggest areas for future improvement [92]. The effectiveness of GANBLR++ demonstrates the value of integrating Bayesian principles into GAN frameworks, providing a robust foundation for capturing complex probabilistic relationships in the data for both numerical and categorical datasets [90].
VIII Future Work
Building upon the insights derived from this study, potential directions for future investigations include:
  1. 1.
    Real vs AI class balance: A predominant challenge in ML tasks is ensuring balanced classes for unbiased predictions. This requirement raises questions about the real-world balance of attack classes, necessitating thorough examination. It is crucial to scrutinize whether achieving balance among ML classes facilitates accurate data synthesis and optimizes ML tasks.
  2. 2.
    Benchmarking the boundary of attack and normal traffic: This research applies generative AI techniques to analyze the distinction between normal and abnormal network traffic rigorously. Through this exploration, we aim to enhance our comprehension of threshold dynamics, significantly contributing to developing more resilient IDS.
  3. 3.
    Fidelity vs Privacy: Enhancing the fidelity and utility of synthetic data often poses privacy risks. Thus, it is imperative to identify an optimized balance between fidelity and privacy, particularly concerning cybersecurity datasets. This endeavor requires careful consideration to safeguard individual privacy while maintaining the data’s utility for cybersecurity purposes.
  4. 4.
    Data dynamic features: Develop and refine feature selection methods targeting data dynamic features in network flows. These techniques should adapt to evolving network environments, identifying the most relevant features for real-time IDS.
  5. 5.
    Adaptive IDS: Design IDS that dynamically adjusts to changes in network traffic patterns. This would involve real-time analysis and feature selection to detect new threats emerging as the network evolves.
  6. 6.
    Loss functions: Conduct a comparative analysis of various generative AI loss functions tailored to the study’s specific context. This will help identify which loss functions are most effective in optimizing the application’s model performance, providing insights into their strengths and limitations in different scenarios.
  7. 7.
    Correlation vs. Dependency: To examine the deeper relationships between variables by exploring logical and conditional dependencies, going beyond the surface-level insights provided by simple correlation measures. This will allow a more critical understanding of how variables interact, particularly in complex datasets where correlations may not fully capture the underlying dependencies.
IX Conclusion
Our comparative study systematically explores the effectiveness of synthetic tabular data generation techniques within cybersecurity network traffic, covering both daily and malicious traffic scenarios. Through a detailed evaluation of several synthetic data generation methods across two significant datasets, we have revealed critical insights into their performances and the inherent trade-offs they entail. SMOTE and Cluster Centroids demonstrated substantial data similarity and class balance but suffer from the drawback of producing redundant, non-novel values. Conversely, CTGAN and TVAE outperformed the other methods, although they do not inherently ensure balanced class distribution. CastGAN emerges as a strong contender by maintaining internal feature dependencies, offering a more balanced approach. The insights from this study provide essential guidance for selecting synthetic data generation techniques based on specific cybersecurity requirements. This research significantly impacts the field by highlighting the strengths and limitations of each method, offering a clear framework for optimizing synthetic data generation in cybersecurity applications. The source code of this experiment is available at our https://github.com/AdenRajput/Comparative_Analysis.git.
X Acknowledgements
The work presented here was partly financed by the European Celtic+ and Swedish Vinnova project ”CISSAN – Collective Intelligence Supported by Security Aware Nodes”.

### v2

@@
-Synthetic Data Generation in Cybersecurity: A Comparative Analysis
+Synthetic Network Traffic Data Generation: A Comparative Study
 Abstract
-Synthetic data generation faces significant challenges in accurately replicating real data, particularly with tabular data, where achieving high fidelity and utility is critical.
-While numerous methods have been developed, the most effective approach for creating high-quality synthetic data for network traffic security remains to be seen.
-This study conducts a comprehensive comparative analysis of non-AI, conventional AI, and generative AI techniques for synthetic tabular data generation using two widely recognized cybersecurity datasets: NSL-KDD and CICIDS-2017.
-Particular emphasis was placed on prominent GAN models for tabular data generation, including CTGAN, CopulaGAN, GANBLR++, and CastGAN.
-The results indicate that GAN-based methods, particularly CTGAN and CopulaGAN, outperform non-AI and conventional AI approaches in terms of fidelity and utility.
-To the best of our knowledge, this research contributes to the field by offering the first comparative evaluation of these methods specifically for cybersecurity network traffic data, filling a critical gap in the literature.
-It also introduces mutual information for feature selection, further enhancing the quality of the generated synthetic data.
-These findings provide valuable guidance for researchers seeking the most suitable synthetic data generation method in cybersecurity applications.
+The generation of synthetic network traffic data is essential for network security testing, machine learning model training, and performance analysis.
+However, existing methods for synthetic data generation differ significantly in their ability to maintain statistical fidelity, utility for classification tasks, and class balance.
+This study presents a comparative analysis of twelve synthetic network traffic data generation methods, encompassing non-AI (statistical), classical AI, and generative AI techniques.
+Using NSL-KDD and CIC-IDS2017 datasets, we evaluate the fidelity, utility, class balance, and scalability of these methods under standardized performance metrics.
+Results demonstrate that GAN-based models, particularly CTGAN and CopulaGAN, achieve superior fidelity and utility, making them ideal for high-quality synthetic data generation.
+Statistical methods such as SMOTE and Cluster Centroid effectively maintain class balance but fail to capture complex traffic structures.
+Meanwhile, diffusion models exhibit computational inefficiencies, limiting their scalability.
+Our findings provide a structured benchmarking framework for selecting the most suitable synthetic data generation techniques for network traffic analysis and cybersecurity applications.
 Index Terms:
-Generative Adversarial Networks (GANs), Synthetic tabular Data, Cybersecurity, NSL-KDD, CICIDS, Mutual information, Feature selection
+Synthetic tabular Data, Network traffic, NSL-KDD, CIC-IDS, Mutual information, Generative Adversarial Networks (GANs), Diffusion models, generative AI
 I Introduction
-In the rapidly evolving field of cybersecurity, IDS plays a crucial role in identifying and mitigating threats within network environments.
-These systems rely heavily on high-quality data to detect patterns indicative of malicious activity [1].
-However, acquiring real-world cybersecurity data poses significant challenges, primarily due to privacy concerns and time constraints [2].
-Researchers have turned to synthetic data generation as a viable alternative to address this.
-Synthetic data, when adequately generated, can mimic the characteristics of real-world data while avoiding the associated ethical and legal issues, making it a valuable resource for training and evaluating IDS [3].
-Despite the growing interest in synthetic data generation, producing synthetic datasets that maintain the utility, fidelity, and data integrity required for practical cybersecurity applications remains a significant challenge [4].
-Utility refers to the ability of synthetic data to support the same analytical tasks as real data [5].
-At the same time, fidelity ensures that the synthetic data accurately reflects the statistical properties of the original data[6].
-Data integrity, on the other hand, focuses on the accuracy and reliability of data, ensuring that it remains unaltered and consistent [7, 8].
-Follow these criteria to avoid synthetic data that is either unrealistic or similar to real data, rendering it less valuable or risky.
-GANs have shown great potential for generating synthetic tabular data for cybersecurity purposes.
-Various GAN architectures like VanillaGAN[9], WGAN[10], and WGAN-GP[11], as well as models specifically designed for tabular data generation such as CTGAN[12], CopulaGAN [13], and GANBLR [14], have been successfully used to create synthetic IDS data [15].
-However, the effectiveness of GANs for generating tabular data specific to cybersecurity, such as network flow data used in IDS, has yet to be thoroughly benchmarked against other tabular synthetic data generation methods: Artificial intelligence (AI) based and non-AI based.
-I-A Synthetic data for cybersecurity
+The increasing reliance on data-driven decision-making in networking and security has intensified the need for high-quality network traffic data.
+This data plays a fundamental role in various applications, including network performance analysis, security research, and the development of machine learning (ML) models [1].
+However, collecting and utilizing real-world network traffic data presents significant challenges.
+Issues such as privacy concerns, limited availability, and the high cost of manual labeling restrict access to large, diverse, and representative datasets [2].
+To overcome these limitations, synthetic network traffic data generation has emerged as a promising alternative [3, 4, 5, 6, 7].
+Synthetic datasets can simulate network traffic patterns while mitigating ethical, legal, and confidentiality concerns, enabling researchers and practitioners to develop, evaluate, and deploy data-driven solutions in a more scalable manner [8].
+Despite this potential, existing methods for synthetic network traffic generation vary widely in their effectiveness, and no comprehensive comparison exists to guide the selection of the most suitable approach for synthetic network traffic data generation.
+I-A The Need for Synthetic Network Traffic Data
+Synthetic network traffic data generation is important for several reasons:
   1. 1.
-    Data privacy: Sharing real-world cybersecurity data involves a risk of sharing sensitive information, which hinders collaboration and research [16].
-Thus, the synthetic data resolves this issue, and valuable data can be shared without compromising privacy [17].
+    Privacy and Data Sensitivity: Due to regulatory restrictions (e.g., GDPR) and commercial confidentiality, real network traffic data is often inaccessible [9].
+Synthetic data provides a privacy-preserving alternative without exposing sensitive user information [10].
   2. 2.
-    Data scarcity: The real-world cyber security data is often imbalanced, with normal traffic making up a higher percentage than attacks or malicious traffic.
-This causes a problem of data scarcity.
-Synthetic data in this scenario can help balance both instances and improve cyber security tasks, such as efficient, unbiased intrusion detection [1].
+    Data Imbalance and Scarcity: Real-world network traffic datasets are highly imbalanced, with benign traffic dominating over malicious activity, which can limit the effectiveness of ML-based models [11].
+Synthetic data can address this by generating balanced datasets for better model training and evaluation.
   3. 3.
-    Diversity of scenarios: Cybersecurity threats are becoming more dynamic and diverse, but real cybersecurity datasets often need more diversity [18].
-In this scenario, synthetic data helps simulate various uncommon attacks and allows for more comprehensive testing of cybersecurity tools [19].
+    Scenario Diversity: Real-world datasets may not cover a broad range of traffic conditions.
+Synthetic data generation techniques allow for the creation of diverse traffic scenarios, including rare or unseen network behaviors, which are crucial for testing robust ML models [12, 13].
   4. 4.
-    Enhancing AI models: When training cyber security models such as Intrusion detection systems (IDS), a large, high-quality dataset is essential for effective detection.
-Therefore, synthetic data plays a crucial role in training these models on a diverse and extensive dataset to enhance the robustness of AI models [20, 21].
-I-B Synthetic data challenges in cybersecurity
+    Scalability and Adaptability: Unlike real data collection, which requires extensive monitoring infrastructure, synthetic data can be generated on demand to match specific analytical or experimental needs, reducing reliance on costly and time-consuming data collection processes [14].
+I-B Challenges in Synthetic Network Traffic Generation
+Despite its advantages, generating high-quality synthetic network traffic data remains challenging:
   1. 1.
-    Data realism: One of the primary challenges in generating synthetic data is ensuring its realism compared to real-world data [22].
-The effectiveness of AI models trained on synthetic data depends on how accurately the data represents real cyber threats and network conditions[23].
-Many generative models are still in the process of achieving data realism by increasing their fidelity and utility to mimic actual trends [24].
+    Ensuring Realism: Many synthetic data generation methods struggle to replicate the statistical and temporal properties of real network traffic, limiting their practical applicability [15].
   2. 2.
-    Lack of variability: Some generative models are explicitly designed to generate a diverse set of network instances and cyber threats.
-However, if synthetic data is not generated with sufficient variability, it may miss capturing the full spectrum of possible attack vectors in cybersecurity [25, 26].
-For instance, the Wasserstein GAN has solved the problem faced by the vanilla GAN of mode collapse, thus ensuring that the generated outcome is not limited to a few instances [10].
+    Variability and Generalization: Some generative models suffer from mode collapse, producing limited traffic variations, which reduces their effectiveness in representing diverse network conditions [16, 17, 18].
   3. 3.
-    Class imbalance: Cybersecurity often suffers from class imbalance, where regular traffic is much larger than cyberattacks [27].
-Synthetic data generation methods address this problem by balancing malicious and benign traffic and attacks within the cybersecurity dataset for more robust intrusion detection [28].
+    Computational Complexity: State-of-the-art generative methods, such as Generative Adversarial Networks (GANs) and diffusion models, require significant computational resources, which can hinder their scalability for real-world deployment [19].
   4. 4.
-    Privacy concerns: Synthetic data attempts to conceal confidential information from real-world sensitive data to address privacy and security concerns [29].
-However, it is not entirely free from these challenges.
-Ensuring that sensitive information cannot be traced back from the synthetic data is crucial and remains a concern.
-  5. 5.
-    Standardization: More standardization is needed for generating, validating, and integrating synthetic data into cybersecurity practices.
-This makes it challenging to ensure the reliability and consistency of the synthetic datasets across different cybersecurity applications [30, 31].
-  6. 6.
-    Perfomace evaluation: For evaluating synthetic data, it is crucial to note that achieving a high similarity score does not guarantee its real-world applicability.
-Similarly, obtaining high accuracy in synthetic cybersecurity datasets for any ML/DL model does not ensure compatibility with real-world datasets [24, 32].
+    Class Imbalance: Many real-world network datasets exhibit severe class imbalances, where attack or rare traffic classes are underrepresented.
+Existing synthetic data generation techniques often fail to address this imbalance effectively, leading to biased models that struggle with anomaly detection and classification [20].
 I-C Motivation
-This research addresses critical challenges in strengthening IDS defenses amidst a rapidly escalating cyber threat landscape.
-IDS are vital for securing networks by detecting and mitigating intrusions, but a lack of high-quality, real-world data for training and refinement hampers their effectiveness.
-Obtaining such data is challenging due to privacy concerns, ethical issues, and evolving network behaviors.
-Synthetic data generation using techniques like GANs offers a promising solution by creating realistic, high-fidelity data that mimics network activities without violating privacy or legal constraints.
-This research is crucial to improving IDS capabilities, enabling them to adapt to modern cyber adversaries’ sophisticated tactics.
-This study aims to enhance IDS robustness and reliability by benchmarking different synthetic data generation methods, contributing to more secure cybersecurity frameworks.
-This research fills a gap in cybersecurity by offering a comprehensive comparison of synthetic data generation methods for enhancing IDS.
-While synthetic data’s potential has been recognized, there is limited clarity on which methods—traditional, AI-based, or GAN-based—provide the highest fidelity and utility for IDS.
-Existing literature has explored various techniques but lacks a focused comparison of their effectiveness in generating network traffic data for IDS.
-This study bridges the gap by systematically evaluating different synthetic data generation techniques, guiding the development of more effective IDS and contributing to more robust, more adaptable cybersecurity defenses.
+Generating synthetic network traffic data plays a critical role in network analysis, security testing, and developing machine learning models.
+While various methods exist for generating synthetic traffic, ranging from non-AI (statistical techniques) to AI (classical and generative) approaches, their effectiveness varies significantly [21, 22].
+Prior studies have explored these methods individually, yet their comparative strengths and weaknesses remain unclear.
+One of the key limitations in existing research is the lack of a systematic evaluation of synthetic data generation techniques across different methodological categories.
+Most studies focus on a single class of methods without directly comparing their ability to generate high-fidelity and utility-preserving data.
+Another challenge lies in handling class imbalance in network traffic datasets.
+Many real-world network environments exhibit highly skewed distributions, where benign traffic significantly outnumbers attack or anomaly instances.
+While synthetic data generation has been proposed as a solution, current methods often fail to maintain class balance without compromising data fidelity [23].
+To address these limitations, this study systematically evaluates and compares multiple synthetic network traffic generation methods across non-AI (statistical) and AI ( classical and generative) approaches.
+By assessing their fidelity, utility, and class-balancing capabilities, this research provides valuable insights into selecting the most suitable synthetic data generation method for different network traffic applications.
 I-D Objectives
-The primary objective of this study is to rigorously evaluate and compare the performance of various synthetic data generation methods for cybersecurity, focusing on their utility, fidelity, data integrity, and ability to handle class imbalance.
-To achieve this, the research addresses the following key questions:
+This study aims to systematically evaluate synthetic network traffic data generation methods by addressing the following research questions:
   1. 1.
-    Which methods produce synthetic data that best preserves the utility for training IDS models?
+    Fidelity and Realism: How well do different synthetic data generation techniques preserve real network traffic’s statistical properties and temporal dynamics?
   2. 2.
-    How well do these methods maintain the original data’s statistical properties (fidelity) and integrity?
+    Utility for Machine Learning Tasks: How effectively can synthetic data generated by different methods improve the performance of ML-based network traffic classification models?
   3. 3.
-    Which techniques effectively balance class distributions in the generated data?
+    Class Balance and Data Diversity: Which methods most effectively address class imbalance in network traffic datasets while maintaining statistical similarity?
   4. 4.
-    Which GAN models are best for producing cyber security network traffic data?
-These questions guide the comprehensive benchmarking of both statistical and AI-based synthetic data generation techniques.
-I-E Contribution
-The contributions of this work are summarized below:
-  1. 1.
-    Comprehensive Comparison of Synthetic Data Generation Methods: This research uniquely compares three prominent synthetic data generation methods—statistical, classical AI, and generative AI—on cybersecurity network traffic datasets (NSLKDD, CICIDS17).
-This provides the first direct comparison of these methods in the cybersecurity field.
-  2. 2.
-    Filling the Knowledge Gap in Cybersecurity Data: The study addresses a gap in the literature by identifying the most suitable synthetic data generation method specifically for network traffic datasets, offering clarity in an area where no previous comparative study exists.
-  3. 3.
-    Practical Guidance for Researchers: By evaluating the performance of these methods, this research offers a clear recommendation on the best approach for generating synthetic cybersecurity data, enabling researchers to make informed decisions without having to test multiple methods.
-  4. 4.
-    Introduction of Mutual Information for Feature Selection: Using Mutual Information for feature selection in synthetic data generation is an innovative contribution, providing a new perspective that may improve feature selection strategies in future cybersecurity studies.
-The source code of this experiment is publicly available on our GitHub Repository  [^1] https://github.com/AdenRajput/Comparative_Analysis.git.
+    Scalability and Computational Efficiency: What are the computational trade-offs associated with each method, and how do they impact scalability for large-scale simulations?
+I-E Contributions
+This study presents a systematic benchmarking of synthetic network traffic generation methods across statistical (non-AI) and AI (classical and generative) AI. Our key contributions are:
+1) Comprehensive Comparative Analysis: We systematically evaluate twelve prominent synthetic data generation methods using standardized metrics, including fidelity, utility for machine learning models, class balance, and computational efficiency.
+This study provides a structured comparison under a unified experimental framework, addressing the lack of direct performance comparisons in prior work.
+2) Empirical Insights into Trade-offs Between Fidelity, Utility, and Computational Cost: Our analysis reveals that generative AI models achieve superior fidelity but require higher computational resources and often struggle with class balance.
+In contrast, statistical methods effectively handle class imbalance but fail to preserve complex traffic patterns.
+These findings provide practical recommendations for selecting the most suitable method based on specific application needs.
+3) Evaluation of Class Imbalance Handling in Synthetic Network Traffic Generation: We analyze how different methods address the issue of class imbalance, a persistent challenge in real-world network traffic datasets.
+Our results highlight the effectiveness of various techniques in generating balanced yet realistic traffic distributions, contributing to the development of more reliable synthetic datasets for machine learning applications.
+The source code for this study is publicly available at https://github.com/AdenRajput/Comparative_Analysis.git.
 I-F Paper Organization
-The remaining sections of this paper are organized as follows:
-  • Section II provides an overview of Synthetic Tabular Data Generation Methods, their usage in tabular data applications in cybersecurity, and evaluation metrics.
-  • Section III details the prominent standard non-AI and AI-based synthetic tabular data generation methods and their mathematical workflow.
-  • Sections IV and V describe the mutual information and evaluation metrics used in this experiment.
-  • Section VI discusses the experimental setup, including the datasets and model configurations.
-  • Section VII presents our experiments’ results, highlighting each method’s performance across two datasets.
-  • Finally, in Sections VIII and IX, we mention the potential areas for future work and development, focusing on critical aspects that require further investigation and improvement, and conclude with key highlights of this study.
+The rest of this paper is structured as follows:
+  • Section II reviews existing synthetic data generation methods and highlights gaps in prior research.
+  • Section III describes the non-AI (statistical) and AI (classical and generative) based synthetic tabular data generation methods.
+  • Section IV introduces the evaluation metrics used to assess fidelity, utility, and class balance.
+  • Section V presents the experimental setup, including datasets, pre-processing steps, and model configurations.
+  • Section VI provides a detailed comparative analysis of results across different methods.
+  • Section VII outlines future research directions and potential improvements in synthetic network traffic generation.
+Figure 1: Methods for generating synthetic tabular data
+TABLE I: Comparison of Synthetic Network Traffic Data Generation Methods
+Category Method Key Features Strengths Weaknesses
+Non-AI (Statistical) Random Oversampling (ROS) [24] Randomly duplicates minority class samples Simple, low computational cost Can lead to overfitting
+Synthetic Minority Oversampling Technique (SMOTE) [25] Generates synthetic samples via interpolation Generalized decision boundaries for classifiers and reduce overfitting Doesn’t consider majority class
+Adaptive Synthetic Sampling (ADASYN) [26] Adaptive sample generation focused on hard-to-learn examples Addresses class imbalance dynamically Can introduce noisy samples
+Cluster Centroids (CC) [27] Reduces majority class by replacing samples with centroids Avoids data duplication May oversimplify and loss key information about data structure
+Gaussian Mixture Models (GMM) [28] Models data distribution probabilistically Captures complex distributions Sensitive to parameter tuning
+AI-Based (Classical + Generative) Bayesian Networks [29] Probabilistic graphical models Captures dependencies between features Requires expert knowledge for structure learning
+Tabular Variational Autoencoder (TVAE) [30] Encode-decoder architecture for learning latent representations Learns complex distributions Computationally expensive
+Tabular Diffusion Model (TabDDPM) [31] Uses iterative denoising to generate data Preserves feature dependencies Computationally intensive
+Conditional Tabular GAN (CTGAN) [32] Conditional generator and mode-specific normalization Handles mixed data types effectively Requires extensive hyperparameter tuning
+CopulaGAN [33] Uses copula transformations in GANs Preserves feature dependencies Less effective in high-dimensional data
+Naive Bayes and Logistic Regression GAN (GANBLR++) [34] Enhances GAN training with Bayesian learning Generates both categorical and numerical data using a Dirichlet Mixture Mode Requires accurate numerical indices
+Cascaded Tabular GAN (CasTGAN) [35] Cascade training mechanism for tabular data Preserves feature dependencies Computationally intensive
 II Related Work
-II-A Synthetic Tabular Data Generation Methods
-Synthetic data is an artificially produced dataset that mimics the statistical features of real-world data without disclosing any sensitive information [3].
-This data type spans various forms, such as tabular data, images, and text.
-Tabular data, organized in a table format with rows and columns, is widely used for ML and DL modeling.
-It is also crucial for cybersecurity applications due to its structured nature.
-It is ideal for representing logs, transaction records, and other important information for detecting and mitigating security threats [16].
-For example, cybersecurity datasets often contain tabular data in IDS, fraud detection, and anomaly detection, enabling precise analysis and decision-making [33, 34].
-Additionally, tabular data’s clear and organized format allows for effectively applying ML and DL algorithms to enhance security measures [22].
-The significance of synthetic data in cybersecurity stems from the challenges associated with accessing high-quality, real-world data in this field.
-Privacy concerns, regulatory limits, and the sensitive nature of cybersecurity information often restrict the availability of data [35].
-By facilitating the generation of realistic datasets free of genuine, sensitive information, synthetic data offers a solution to these challenges [36].
-This advancement is critical for pushing forward research, enhancing security models, and executing thorough threat analyses.
-II-B Applications in Cybersecurity
-Synthetic data finds multiple applications in cybersecurity, improving different facets of security management and threat response:
-  1. 1.
-    Intrusion Detection Systems (IDS): By supplying varied and representative datasets encompassing a range of attack patterns and behaviors, synthetic data aids in the training and testing of IDS, enhancing their detection accuracy and resilience [37].
-Research has shown that employing synthetic data techniques on datasets like UNSW-NB15 and CICIDS-17 has significantly improved the performance of deep learning models used in IDS [38].
-  2. 2.
-    Anomaly Detection: Synthetic data plays a role in simulating typical and atypical behaviors within networks, thereby facilitating the training of systems to detect deviations indicative of potential security incidents [39].
-Synthetic data has been used to preserve privacy and detect anomalies in the KDD dataset.
-Its effectiveness, measured in supervised, semi-supervised, and unsupervised anomaly detection, showed an accuracy of 99.69%, similar to models trained on original data [40].
-  3. 3.
-    Security Training and Simulation: The value of synthetic data extends to crafting realistic scenarios for security training exercises and simulations.
-This enables security teams to hone their responses to varied cyber threats without compromising real data [41].
-A deep Convolutional Generative Adversarial Network (DCGAN) model was used to generate synthetic cyberattack data and realistic attack simulations to train deep learning models for better detection.
-The model with those simulated attacks achieved an accuracy of 99.69% on the KDD dataset and 97.93% on the CICIDS-17 dataset [42].
-In discussing the significance and uses of synthetic tabular data in cybersecurity, we will focus on the methods used to generate such data effectively.
-The later sections will delve into the specifics and their architecture.
-They will also provide an experimental analysis of the prominent standard (non-AI) and AI-based methods for creating synthetic tabular data for cybersecurity IDS.
-II-C Metrics for Evaluation
-  1. 1.
-    Fidelity: Fidelity measures how well the synthetic data replicates the statistical properties of the original real-world data [43].
-Therefore, the single most significant metric for evaluating synthetic tabular data generally involves assessing the data’s underlying distribution [44].
-This approach is based on the definition of data synthesis to capture the probability distribution function of the original data, thereby creating synthetic data that mirrors these characteristics [45].
-However, the evaluation process can be subjective, based on the data production’s objective.
-As a result, researchers employ various methods to assess synthetic tabular data, considering its structure markedly differs from that of images.
-    Unlike synthetic tabular data, the validation of synthetic image-based output is often visual.
-If the model yields realistic images, it’s deemed successful.
-Images comprise spatial information, shape, texture, objects, and color [46].
-Nonetheless, synthetic data is confined to images and encompasses tabular data [47], characterized by its complexity, multiple interconnected features, hidden patterns, and context-dependent information [48].
-Hence, evaluating synthetic tabular data demands a more structured approach.
-  2. 2.
-    Utility: Utility measures the usefulness of synthetic data for the same tasks or analyses as the original data [49].
-To evaluate the quality of the synthetic data, key performance metrics such as accuracy, F1 score, precision, and recall were compared between models trained and tested on real data versus those involving synthetic data[50, 15, 51].
-Specifically, the following scenarios were used to rank performance:
-    TRTR (Train and Test on Real Data): Baseline performance using real data.
-TSTR (Train on Synthetic, Test on Real): Measures how well models trained on synthetic data generalize to real data.
-TRTS (Train on Real, Test on Synthetic): Assesses the consistency of synthetic data in replicating real data behavior.
-TSTS (Train and Test on Synthetic Data): Evaluates synthetic data’s internal consistency and utility within a self-contained model [52, 53].
-By comparing these four scenarios, the utility of the synthetic data can be ranked, providing insight into how well it replicates the practical use of real data [54, 55, 56].
-  3. 3.
-    Universal metrices: Recent research has examined synthetic tabular data via diverse statistical and ML methods [57, 58, 59].
-Nevertheless, these approaches could be more precise regarding suitable evaluation measures.
-Hence, there’s been a move to define universal metrics to boost universality and interpretability.
-One such metric is TabSynDex, offering a score between 0 and 1 to gauge the synthetic data’s closeness to real data [60].
-However, this metric still needs to improve in selecting the proper evaluation measures.
-For example, there needs confirmation that TabSynDex’s five main components adequately represent the synthetic data’s completeness or facilitate more accessible validation of generated tabular data.
-A recent SCH Yang et al.
-proposal introduces a structured evaluation methodology focused on comparing the distributions of synthetic and real data [44].
-This study is pivotal as it aligns with generative AI’s core principle: replicating the probability distribution of the original dataset.
-Therefore, by considering the recent research in the domain, this study employs prominent evaluation metrics such as fidelity (statistical similarity), utility, and class balance in its evaluation section.
-III Synthetic Data Generation Techniques
-Figure 1: Methods for generating synthetic tabular data
-III-A Standard Methods (non-AI)
-  1. 1.
-    Random Oversampling (ROS) is a widely used technique to address class imbalance, a common issue in datasets, especially in domains like cybersecurity where certain types of data, such as network intrusions, are rare compared to normal traffic.
-ROS works by duplicating samples from the minority class to balance the class distribution, thereby preventing ML models from becoming biased toward the majority class [61].
-This technique is particularly relevant when the non-complex small dataset lacks sufficient examples of minority classes.
-    In synthetic tabular data generation, ROS is often employed to artificially increase the number of minority class samples, helping models better learn the characteristics of these rare events.
-However, its application in cybersecurity data, which often involves complex relationships between variables, may be limited if used alone, as it does not generate new, diverse data points [62].
-    The basic procedure of ROS is straightforward:
-    1. (a)
-      Identify the Minority Class: Determine which class has fewer samples.
-    2. (b)
-      Duplicate Samples: Randomly select samples from the minority class and duplicate them until the class distribution is balanced.
-    3. (c)
-      Augmented Dataset: The result is a dataset with a balanced class distribution, where the minority class is equally represented as the majority class
-    This process does not alter the duplicated samples, meaning the new data points are exact copies of the original samples.
-  2. 2.
-    Synthetic Minority Oversampling Technique (SMOTE) is an advanced over-sampling method designed to address the limitations of ROS by generating synthetic examples rather than merely duplicating existing ones.
-Introduced by Chawla et al.
-in 2002, SMOTE aims to balance the class distribution in datasets by creating new instances of the minority class, which helps improve the performance of ML models, particularly in scenarios where the class imbalance is significant [63].
-Unlike traditional over-sampling methods, SMOTE synthesizes new samples by interpolating between existing minority class samples, leading to more generalized decision boundaries for classifiers [64].
-    SMOTE is particularly useful in domains such as cybersecurity, where the occurrence of attacks (minority class) is much rarer than normal behavior (majority class).
-The technique is well-suited for situations where the minority class instances are spread sparsely across the feature space, and duplicating existing samples (as done in ROS) could lead to overfitting [65, 66].
-    The SMOTE algorithm operates through the following steps:
-    1. (a)
-      Minority Class: For each minority class instance, there is a need to select k-nearest neighbors, e.g., k = 5
-    2. (b)
-      Synthetic Samples: One of the k-nearest neighbors is randomly chosen to generate a synthetic data point.
-The synthetic data point is created by interpolating between the feature vector of the original data point and that of the selected neighbor.
-This interpolation involves calculating the difference between the feature vector of the selected neighbor and the original data point.
-Then, this difference is multiplied by a random number between 0 and 1, and the resulting value is added to the feature vector of the original data point.
-    3. (c)
-      Augmented Dataset: The result is a dataset augmented with synthetic samples that exhibit variations within the minority class, thereby mitigating the risk of overfitting and improving the model’s ability to generalize.
-  3. 3.
-    Adaptive Synthetic Sampling (ADASYN) is an oversampling method similar to ROS and SMOTE.
-However, it introduces a more adaptive approach to generating synthetic data.
-It was introduced by He et al.
-in 2008 and focuses on the harder-to-learn examples from the minority class.
-ADASYN improves on SMOTE by adapting the sampling procedure based on the learning difficulty of each minority class sample.
-Data points that are harder to classify receive more synthetic neighbors; the rest receive fewer neighbors.
-[67].
-    ADASYN is particularly useful in domains such as cybersecurity, where the rarity of specific attack types (minority class) can cause models to perform poorly [68].
-By adaptively generating synthetic samples based on the distribution of the minority class, ADASYN helps models learn the decision boundary more effectively, leading to better generalization and improved detection of minority class instances [69].
-    ADASYN operates through the following steps:
-    1. (a)
-      Identify Difficult Minority Class: ADASYN computes the local density of each minority class by analyzing how many of its nearest neighbors belong to the majority class.
-Data points surrounded by majority-class neighbors are considered harder to classify.
-    2. (b)
-      Weight Assignment: Weights are assigned to each minority instance/data point based on difficulty level.
-Harder-to-classify points receive a higher weight, implying that more synthetic samples will be generated.
-    3. (c)
-      Synthetic Samples: For each minority instance, ADASYN interpolates between the instances and one of its nearest minority class neighbors.
-Synthetic samples are created along the line connecting the two instances.
-    4. (d)
-      Data Balancing: ADASYN adds the generated samples to the original ones, focusing on the problematic region to reduce the class imbalance and improve the model training and learning.
-  4. 4.
-    Cluster Centroids is an under-sampling technique designed to address class imbalance by reducing the number of majority class instances.
-Unlike traditional under-sampling methods, which randomly remove majority class samples, Cluster Centroids utilizes a clustering algorithm to identify representative centroids within the majority class.
-The technique was popularized to preserve the majority class’s distribution and structure while reducing its size, thereby balancing the dataset without losing critical information [70, 71].
-This approach is beneficial in scenarios where the majority class is substantial, and random under-sampling could significantly lose valuable data.
-    The concept of cluster centroids is applicable in various domains, including cybersecurity.
-In highly imbalanced datasets, where normal traffic significantly outweighs the attack data, cluster centroids help by grouping similar attack (minority class) samples into clusters and generating synthetic data points at the centroid of each cluster.
-This helps represent the attack patterns more effectively and exposes ML models to diverse attacks for robust anomaly detection [72].
-    Cluster Centroids operates through the following steps:
-    1. (a)
-      Minority Class Clustering: Minority class data is grouped into clusters using a clustering algorithm, e.g., K-Means clustering. Each cluster contains similar data points from the minority class.
-    2. (b)
-      Centroid Calculation: For each cluster, the algorithm computes the centroids, the central points representing the average position of the data points within a particular cluster in the feature space.
-    3. (c)
-      Synthetic Data: Synthetic samples are generated at or around the center of each minority cluster.
-This ensures that the generated data is representative of the overall distribution of the minority class.
-    4. (d)
-      Balanced Dataset: Synthetic data is added to the original data to balance class distributions, improving the AI model’s ability to learn from minority classes for higher efficiency.
-    Figure 2: Bayesian Network illustrating some key feature dependencies in the NSL-KDD dataset
-  5. 5.
-    Gaussian Mixture Model (GMM) is a probabilistic model used to represent the presence of sub-populations within an overall population without requiring that an instance belong exclusively to a single sub-population.
-GMMs are particularly useful for modeling data that exhibit multiple underlying distributions, assuming that all data points are generated from a mixture of several Gaussian distributions with unknown parameters [73].
-Each Gaussian component within a GMM represents a cluster characterized by its mean, variance, and mixing coefficient.
-    In the context of synthetic tabular data generation, GMMs approximate the data distribution, making them highly applicable for generating new samples that preserve the statistical properties of the original dataset.
-This is especially useful in cybersecurity, where modeling the intricate patterns of network traffic or attack vectors requires flexible and accurate representation of complex, multimodal distributions [74].
-GMMs provide a robust method for capturing these patterns, allowing for generating synthetic data that closely mimic real-world scenarios [75].
-    GMM operates through the following steps:
-    1. (a)
-      Fitting GMM: GMM fits the original data by assuming it comes from the mixture of multiple Gaussian distributions.
-It estimates these Gaussian components’ parameters (mean, covariance, and weights) using the Expectation-Maximization (EM) algorithm.
-    2. (b)
-      Estimate Gaussian Parameters: The model assigns data points to different Gaussian components based on their likelihood and calculates the mean (center), covariance (spread), and mixture weight (probability and skewness) for each Gaussian distribution.
-    3. (c)
-      Sampling: After training the model, synthetic data is generated by randomly selecting a Gaussian component (based on its weight) and then sampling a point from the selected Gaussian using its parametric values.
-    4. (d)
-      Data Generation: Repeating the sampling process multiple times generates synthetic data that mimics the structure and underlying distribution of the original data.
-Thus, the generated data points from all components are combined to form the final synthetic dataset.
-[76].
-Figure 3: Variational Autoencoder Structure
-    III-B AI based Methods
-  1. 1.
-    Baysian Network (BN) are probabilistic graphical models that represent a set of variables and their conditional dependencies via a directed acyclic graph (DAG) [77].
-Bayesian Networks are particularly effective for modeling complex relationships in data where uncertainty and causality play critical roles [78, 79].
-In figure 2, this BN illustrates that target is a central feature, influenced by several other variables, including service_http, service_private, flag_SF, logged_in, and flad_S0.
-This suggests that target plays a crucial role in determining the system’s behavior, as it is obvious by its binary data type of normal or attack.
-The BN also shows key conditional dependencies, such as flag_SF being dependent on logged_in and service_domain_u, while dst_host_serror_rate influences flad_S0.
-Some features, like service_private and service_http, directly impact target but do not further influence other variables.
-Overall, this network emphasizes the significant joint dependencies between these system features, highlighting target as a critical node and capturing the complex relationships between the various network-related parameters.
-    In synthetic data generation, BNs generate data that preserves the inherent dependencies and conditional probabilities observed in the original dataset.
-This is particularly valuable in cybersecurity, where understanding the intricate dependencies between various network events, user behaviors, or attack patterns is essential for generating realistic synthetic datasets.
-By capturing these relationships, BNs allow for the generation of synthetic data that not only mimics the distribution of the original data but also maintains the underlying causal structures, making it highly applicable for scenarios requiring interpretable and robust synthetic data [14].
-    The process of generating synthetic data using Bayesian Networks involves the following steps [80]:
-    1. (a)
-      Structure Learning: The first step is to determine the structure of the Bayesian Network, which involves identifying the dependencies between variables.
-This can be done using expert knowledge or algorithms that learn the structure from data, such as hill climbing or score-based methods [81].
-Each node in the network corresponds to a variable, and directed edges between nodes represent conditional dependencies.
-    2. (b)
-      Parameter Learning: After defining the structure, the next step is to estimate the conditional probability distributions (CPDs) for each variable in the network, given its parent node.
-This joint probability information is learned from the original dataset.
-    3. (c)
-      Sampling: After learning the structure and parameters (CPDs), synthetic data is generated by sampling from the Bayesian network.
-First, the root node is sampled from its marginal distribution, and then the child nodes are sampled based on their conditional probabilities.
-This process continues throughout the network until all variables are sampled.
-    4. (d)
-      Data Generation: The sampling process is repeated for the required number of synthetic data points.
-Each sample is a synthesized record for the dataset.
-This process ensures that the generated data depicts the dependencies learned from the original dataset.
-  2. 2.
-    Tabular Variational Autoencoder (TVAE) is a generative model designed to handle the complexities of tabular data, often including a mix of continuous and categorical variables.
-TVAE has the base idea of autoencoders that consists of two parts: an encoder network that learns to compress high-dimensional data into a low-dimensional, latent spacial representation (the code) and a decoder network that decompresses the compressed representation into the original domain as depicted in Figure 3 [82].
-Thus, it extends the traditional Variational Autoencoder (VAE) framework by adapting it to the specific needs of tabular data, ensuring that the generated synthetic data accurately reflects the structure and distribution of the original dataset [83].
-    Mathematical process of TVAE for Synthetic data generation:
-    1. (a)
-      Encoder-Decoder Architecture: TVAE uses an encoder-decoder structure where the encoder $q_{\phi}(z|x)$ maps the input data $x$ to a latent space $z$, and the decoder $p_{\theta}(x|z)$ reconstructs the data from this latent representation.
-The encoder outputs the parameters of a Gaussian distribution in the latent space:
-      $q_{\phi}(z|x)=\mathcal{N}(z;\mu_{\phi}(x),\sigma_{\phi}(x)^{2}I)$ (1)
-      Where $\mathcal{N}(z;\mu_{\phi}(x),\sigma_{\phi}(x)^{2}I)$ represents a multivariate normal (Gaussian) distribution where the latent variable $z$ is assumed to follow a Gaussian distribution with mean $\mu_{\phi}(x)$ and covariance matrix $\sigma_{\phi}(x)^{2}I$.
-The two parameters, $\mu_{\phi}(x)$ and $\sigma_{\phi}(x)$, are functions of the input $x$ and are learned by the encoder network.
-More precisely, $\mu_{\phi}(x)$ is the mean vector of the Gaussian distribution, parameterized by the input $x$.
-This vector represents the expected value of the latent variable $z$ for a given input $x$, determining the center of the distribution, and $\sigma_{\phi}(x)^{2}$ is the variance of the latent variable $z$, also parameterized by the input $x$.
-This variance determines how much the latent variable $z$ can deviate from its mean $\mu_{\phi}(x)$, controlling the spread of the distribution.
-Specifically, $\sigma_{\phi}(x)^{2}$ is the element-wise square of the standard deviation $\sigma_{\phi}(x)$.
-Plus, $I$ is the identity matrix used to represent that the latent variables are assumed to be independent, meaning there is no covariance between different latent dimensions.
-The use of $I$ ensures that the covariance matrix is diagonal, implying that each latent variable has its independent variance, controlled by $\sigma_{\phi}(x)^{2}$.
-    2. (b)
-      Variational Objective: The objective of training TVAE is to maximize the Evidence Lower Bound (ELBO):
-      $\begin{split}\mathcal{L}(\theta,\phi;x)=&\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-\\ &D_{KL}(q_{\phi}(z|x)||p(z))\end{split}$ (2)
-      Here, $\mathcal{L}(\theta,\phi;x)$ is the objective function (ELBO) to maximize during the training of TVAE.
-It measures how well the model reconstructs data and regularizes the latent space, and $\log p_{\theta}(x|z)$ represents the reconstruction loss, ensuring that the generated data $\hat{x}$ resembles the input data $x$.
-$D_{KL}$ is the Kullback-Leibler divergence that regularizes the latent space to follow a prior distribution $p(z)$, typically a standard normal distribution $\mathcal{N}(0,I)$.
-Plus, $\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]$ is the expected log-likelihood term, representing the reconstruction loss.
-It ensures that the generated data $\hat{x}$ resembles the input data $x$.
-    3. (c)
-      Latent Variable Modeling: The latent space $z$ is sampled from the Gaussian distribution parameterized by $\mu_{\phi}(x)$ and $\sigma_{\phi}(x)$.
-This latent space effectively captures the dependencies and distribution of the original data.
-    4. (d)
-      Data Generation: To generate synthetic data, samples are drawn from the latent space $z$, and these samples are passed through the decoder $p_{\theta}(x|z)$ to produce synthetic data points $\hat{x}$.
-The decoder ensures that the synthetic data preserves the statistical characteristics of the original tabular dataset.
-    5. (e)
-      Categorical and Continuous Data Handling: TVAE incorporates specific mechanisms to handle the peculiarities of tabular data.
-Standard Gaussian distributions are used for continuous variables, while TVAE often employs a Gumbel-Softmax distribution for categorical variables to allow gradient-based optimization.
-    Figure 4: Diffusion Model Flow
-  3. 3.
-    Diffusion Models, particularly Denoising Diffusion Probabilistic Models (DDPM), have emerged as a powerful class of generative models [84, 85].
-Originally developed for continuous data types such as images, these models have been adapted to handle the complexities of tabular data, which often includes a mix of categorical and numerical variables.
-The application of diffusion models to tabular data is exemplified by the Tabular Diffusion Model (TabDDPM), which has shown state-of-the-art performance in generating high-quality synthetic tabular data [86].
-    Mathematical process of diffusion models for tabular data generation:
-    1. (a)
-      Forward and Reverse Diffusion Processes: The diffusion model begins with a forward process that gradually adds Gaussian noise to the data.
-Given a data point $x_{0}$, the forward process generates a sequence of increasingly noisy versions $x_{1},x_{2},\dots,x_{T}$, where each step adds a small amount of noise (Figure 4):
-      $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (3)
-      Here, $\beta_{t}$ is a variance schedule that controls the amount of noise added at each step $t$.
-The term $\sqrt{1-\beta_{t}}$ represents the decay factor, which ensures that the noise doesn’t overwhelm the signal, and $I$ is the identity matrix indicating that the noise is isotropic.
-      The reverse process aims to generate synthetic data by denoising a sample starting from pure noise.
-The reverse distribution $p_{\theta}(x_{t-1}|x_{t})$ is parameterized by a neural network that predicts the mean and variance at each step:
-      $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{\theta}^{2}I)$ (4)
-      Here, $\mu_{\theta}(x_{t},t)$ is the predicted mean and $\sigma_{\theta}^{2}$ is the predicted variance at time step $t$, both parameterized by the neural network $\theta$.
-The network parameters $\theta$ are optimized to minimize the Kullback-Leibler divergence between the forward and reverse processes, effectively learning to denoise the data step by step.
-    2. (b)
-      Loss Function: The training objective for diffusion models in tabular data is typically the variational lower bound (VLB), which can be simplified to the mean-squared error (MSE) between the predicted noise and the actual noise:
-      $\mathcal{L}(\theta)=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right]$ (5)
-      Here, $\epsilon_{\theta}(x_{t},t)$ is the predicted noise by the model at time step $t$, where $x_{t}$ represents the noisy data and $\epsilon_{\theta}$ is parameterized by the neural network $\theta$.
-The term $\epsilon$ represents the noise added to the data during the forward process.
-The model minimizes the difference between $\epsilon$ and $\epsilon_{\theta}(x_{t},t)$, effectively learning to reverse the noise and reconstruct the data.
-      The expectation $\mathbb{E}_{t,x_{0},\epsilon}$ denotes the averaging over all time steps $t$, initial data $x_{0}$, and noise samples $\epsilon$ applied during the forward process.
-    3. (c)
-      Handling Categorical and Numerical Features: Diffusion models for tabular data, such as TabDDPM, extend the basic framework to handle mixed data types.
-Categorical variables are typically processed using multinomial diffusion, where the data is represented as one-hot encoded vectors, and the noise is added in a way that respects the unconditional nature of the data:
-      $q(x_{t}|x_{t-1})=\text{Cat}(x_{t};(1-\beta_{t})x_{t-1}+\beta_{t}/K)$ (6)
-      Here, $\text{Cat}(x_{t};\cdot)$ represents the categorical distribution, which ensures that the synthetic data adheres to the unconditional nature of the variable.
-The term $\beta_{t}$ is the variance schedule controlling the amount of noise added, while $K$ represents the number of categories.
-$(1-\beta_{t})x_{t-1}+\beta_{t}/K$ maintains the correct distribution of categories by adjusting the probability mass of each category at every time step $t$.
-      Numerical variables are handled using standard Gaussian diffusion processes.
-The model learns to denoise both categorical and numerical features in a unified framework, ensuring that the generated synthetic data maintains the relationships and distributions of the original data.
-    4. (d)
-      Synthetic Data Generation: After training, the model can generate synthetic tabular data by sampling from the learned distribution.
-The reverse process generates a sequence of less noisy samples from a sample of pure noise until the final synthetic data point is produced.
-This approach allows the model to create realistic tabular data that captures individual feature distributions and the dependencies between features.
-    Figure 5: Original GAN structure by Goodfellow [9]
-  4. 4.
-    Generative Adversarial Networks (GAN)
-Conditional Tabular GAN (CTGAN)
-CTGAN is a generative model designed explicitly for tabular data, which is often characterized by the presence of mixed data types, including both categorical and continuous variables [12].
-CTGAN extends the traditional GAN framework (Figure 5) by incorporating techniques that are well-suited for the unique challenges of tabular data generation, such as handling imbalanced datasets and preserving the relationships between variables [9, 3].
-Mathematical process of CTGAN for synthetic data generation:
-  1. 1.
-    Generator and Discriminator Networks: CTGAN uses a generator network $G(z,c)$ that takes as input a noise vector $z$ sampled from a standard normal distribution and a conditional vector $c$ that encodes the categorical features.
-The generator produces synthetic data samples that mimic the real data distribution.
-The discriminator network $D(x,c)$ is then tasked with distinguishing between real data samples $x$ and synthetic samples generated by $G(z,c)$ while also considering the conditional vector $c$.
-  2. 2.
-    Mode-Specific Normalization for Continuous Data: To handle the diverse distribution of continuous features in tabular data, CTGAN introduces a mode-specific normalization technique.
-This approach transforms each continuous variable by centering it around a randomly sampled value from its distribution and normalizing it based on the empirical standard deviation of the data within that mode.
-This process ensures that the generator can more effectively model continuous data distributions:
-    $x_{normalized}=\frac{x-x_{sample}}{\sigma(x_{mode})}$ (7)
-    where $x$ is the original continuous variable, $x_{sample}$ is a value randomly sampled from the distribution of $x$, and $\sigma(x_{mode})$ is the empirical standard deviation of $x$ within its mode.
-  3. 3.
-    Conditional Vector Construction: CTGAN utilizes a conditional vector $c$ to guide the generation process, particularly for categorical variables.
-This vector is one-hot encoded, allowing the generator to focus on producing samples that match the specific conditions defined by $c$.
-The conditional vector is also integrated into the discriminator’s training process, helping it better differentiate between natural and synthetic data based on the specified conditions.
-  4. 4.
-    Training Objective: The training process of CTGAN follows the standard GAN framework, where the generator and discriminator are trained in a min-max game.
-The generator aims to minimize the following loss function to fool the discriminator, while the discriminator tries to maximize its ability to distinguish real from synthetic data:
-    $\begin{split}\min_{G}\max_{D}&\mathbb{E}_{x\sim p_{data}}[\log D(x,c)]+\\ &\mathbb{E}_{z\sim p_{z}}[\log(1-D(G(z,c),c))]\end{split}$ (8)
-    Where $G$ represents the generator, $D$ represents the discriminator, $x$ is a data sample drawn from the real data distribution $p_{data}$, $z$ is a random noise vector drawn from the prior distribution $p_{z}$, $c$ denotes any additional conditional information, $G(z,c)$ does the generator generate the synthetic data, and $D(x,c)$ and $D(G(z,c),c)$ are the discriminator’s outputs for real and synthetic data, respectively.
-  5. 5.
-    Data Sampling and Imbalance Handling: CTGAN employs a training-by-sampling technique to address class imbalance in tabular data.
-The data points are sampled during training to ensure underrepresented classes are adequately represented in the training batches.
-This approach mitigates the impact of imbalanced datasets on the quality of the generated data.
-Copula GAN
-CopulaGAN is a generative model that combines the strengths of copula-based statistical models with the flexibility of GANs to generate synthetic tabular data [87].
-Copulas are mathematical tools that describe the dependence structure between random variables, enabling the modeling of complex dependencies separately from the marginal distributions [88].
-By integrating copulas into the GAN framework, Copula GAN effectively captures the intricate dependencies in tabular data, making it particularly suitable for scenarios where preserving the relationships between variables is crucial [89].
-The mathematical process of Copula-GAN for synthetic data generation:
-  1. 1.
-    Marginal Distribution Estimation: The first step in Copula GAN involves estimating the marginal distributions of each variable in the tabular dataset.
-This can be achieved using kernel density estimation for continuous variables or empirical distributions for categorical variables.
-The goal is to transform the original data into a uniform distribution on the interval $[0,1]$ using the cumulative distribution function (CDF) of each variable:
-    $u_{i}=F_{i}(x_{i})\quad\text{for }i=1,\dots,d$ (9)
-    where $u_{i}$ is the transformed variable, and $F_{i}(x_{i})$ is the CDF of the $i$-th variable.
-  2. 2.
-    Copula Function for Dependency Modeling: Copula GAN models the dependencies between the variables using a copula function once the data is transformed into the copula space.
-The copula function $C$ captures the joint distribution of the transformed variables:
-    $C(u_{1},\dots,u_{d})=\mathbb{P}(U_{1}\leq u_{1},\dots,U_{d}\leq u_{d})$ (10)
-    where $C$ denotes the copula function, $u_{1},\dots,u_{d}$ are the values of the copula-transformed variables, and $U_{1},\dots,U_{d}$ are the random variables corresponding to these transformed values.
-The expression $\mathbb{P}(U_{1}\leq u_{1},\dots,U_{d}\leq u_{d})$ represents the joint probability that $U_{1}$ is less than or equal to $u_{1}$, $U_{2}$ is less than or equal to $u_{2}$, and so on up to $U_{d}$ and $u_{d}$.
-    The copula function separates the dependency structure from the marginal distributions, which is advantageous for generating synthetic data that preserves the original data’s relationships.
-  3. 3.
-    Generator and Discriminator Networks: In Copula GAN, the generator $G(z)$ takes a noise vector $z$ and generates samples in the copula space, aiming to mimic the dependency structure captured by the copula function.
-The discriminator $D(u)$ evaluates whether a given sample from the copula space is real or generated, guiding the generator to produce more realistic samples:
-    $\begin{split}\min_{G}\max_{D}&\mathbb{E}_{u\sim p_{data}}[\log D(u)]+\\ &\mathbb{E}_{z\sim p_{z}}[\log(1-D(G(z)))]\end{split}$ (11)
-  4. 4.
-    Inverse Transformation and Synthetic Data Generation: After training, the generator produces samples in the copula space.
-These samples are then transformed back to the original data space using the inverse of the marginal CDFs:
-    $x_{i}=F_{i}^{-1}(u_{i})\quad\text{for }i=1,\dots,d$ (12)
-    Where $x_{i}$ represents the $i$-th component of the synthetic data in the original data space, $u_{i}$ is the $i$-th element in the sample in the copula space, and $F_{i}^{-1}$ denotes the inverse of the marginal CDF of the $i$-th variable.
-This inverse transformation ensures that the generated data points follow the original marginal distributions while maintaining the dependencies modeled by the copula.
-Naive Bayes and Logistic Regression GAN (GAN- BLR++
-GANBLR++ is an advanced generative model designed to address the limitations of traditional GANs in generating synthetic tabular data, mainly where complex dependencies and mixed data types (categorical and numerical) are involved.
-Unlike its predecessor, GANBLR, which primarily focused on modeling categorical data using restricted Bayesian Networks, GANBLR++ introduces significant enhancements that allow it to generate both categorical and numerical data with greater accuracy [14, 90].
-This model integrates unrestricted Bayesian Networks and introduces a Dirichlet Mixture Model (DMM) for numerical data generation, enabling it to capture the full range of dependencies in real-world tabular datasets [91].
-GANBLR++ is particularly useful in domains where the integrity of the generated data is critical, such as cybersecurity, which offers a robust solution that balances flexibility, accuracy, and computational efficiency.
-The mathematical process of GANBLR++ for synthetic data generation:
-  1. 1.
-    Incorporation of Unrestricted Bayesian Networks: GANBLR++ leverages unrestricted Bayesian Networks to model complex dependencies between variables.
-The network structure is learned through a scoring function that optimizes the log-likelihood of the data while including a regularization term to avoid overfitting.
-This allows the model to represent more intricate relationships between variables than the restricted networks used in GANBLR.
-    $\text{Score}(G:D_{\text{data}})=\text{LL}(G:D_{\text{data}})-\lambda\cdot|G|$ (13)
-    where $\text{Score}(G:D_{data})$ is the score of the network structure $G$ given the data $D_{data}$, $\text{LL}(G:D_{data})$ denotes the log-likelihood of the data under the network $G$, $\lambda$ is the regularization parameter, and $|G|$ represents the complexity of the network $G$.
-The regularization term $\lambda\cdot|G|$ helps prevent overfitting by penalizing more complex network structures.
-  2. 2.
-    Dirichlet Mixture Model for Numerical Data: To handle numerical data effectively, GANBLR++ uses a Dirichlet Mixture Model (DMM).
-The DMM identifies modes within the numerical data distribution and samples from these modes during data generation, ensuring that the generated values accurately reflect the underlying distribution of the original data.
-    $x_{i}\sim\text{N}(\mu_{c},\sigma_{c})$ (14)
-    where $x_{i}$ is the $i$-th sample from the numerical data, $\text{N}(\mu_{c},\sigma_{c})$ denotes a Gaussian distribution with mean $\mu_{c}$ and standard deviation $\sigma_{c}$ for a particular mode $c$.
-The parameters $\mu_{c}$ and $\sigma_{c}$ represent the mean and standard deviation of the Gaussian distribution associated with the mode $c$.
-  3. 3.
-    Mode-Specific Sampling: GANBLR++ further refines the generation of numerical data through mode-specific sampling, which ensures that the generated data points not only match the distributional characteristics of the original data but also maintain the necessary granularity.
-    $\begin{split}&x_{i}\sim\text{N}(\mu_{c},\sigma_{c})\quad\text{subject to }\\ &\alpha_{c}<\frac{x_{i}-\mu_{c}}{\sigma_{c}}\leq\alpha_{c}+1\end{split}$ (15)
-    where $x_{i}$ is the $i$-th sample from the numerical data, $\text{N}(\mu_{c},\sigma_{c})$ denotes a Gaussian distribution with mean $\mu_{c}$ and standard deviation $\sigma_{c}$ for a particular mode $c$, and $\alpha_{c}$ is a mode-specific threshold parameter.
-The constraint $\alpha_{c}<\frac{x_{i}-\mu_{c}}{\sigma_{c}}\leq\alpha_{c}+1$ ensures that $x_{i}$ falls within a specific range relative to the mode $c$, maintaining the granularity of the distribution.
-Cascaded Tabular GAN (CasTGAN)
-CasTGAN is a novel framework designed for generating realistic tabular data [92].
-Traditional GAN models often struggle with capturing the complex dependencies between features in tabular datasets, leading to synthetic data that fails to preserve the validity and interdependencies present in the original data [93].
-CasTGAN addresses this limitation by employing a cascaded architecture where each generator in the sequence is responsible for synthesizing a single feature of the dataset.
-This architecture allows the model to better capture the intricate relationships between features, ensuring that the generated data is not only statistically similar to the real data but also semantically valid.
-The primary motivation behind CasTGAN is to enhance the generation of high-quality synthetic tabular data, which is crucial in scenarios where data privacy is a concern and real data cannot be freely shared or used.
-The mathematical framework of CasTGAN can be broken down into the following steps:
-  1. 1.
-    Cascaded Generators: The model is composed of multiple generators $G_{1},G_{2},\dots,G_{M}$, each responsible for generating one feature of the dataset.
-Each generator $G_{i}$ receives inputs from a noise vector $z$ and the outputs of the previous generator $G_{i-1}$.
-  2. 2.
-    Auxiliary Learners: For each generator $G_{i}$, there is an auxiliary learner $AL_{i}$ that computes a loss specific to the feature generated by $G_{i}$.
-The auxiliary loss guides the generator toward producing more accurate and valid feature values.
-  3. 3.
-    Discriminator: A single discriminator $D$ evaluates the final synthetic dataset, distinguishing between real and generated data.
-The discriminator is trained on the output of the final generator $G_{M}$, and its loss is propagated back through the cascade to improve the performance of all generators.
-  4. 4.
-    Training Objective: The training of CASTGAN involves minimizing the Wasserstein distance with a gradient penalty between the real and generated data distributions.
-The value function for the discriminator and the generators is given by:
-    $\begin{split}&\min_{G_{\rightarrow}}\max_{D}V(D)=\mathbb{E}_{x\sim p_{data}}[D(x)]\\ &-\mathbb{E}_{z\sim p_{z}}[D(G_{\rightarrow}(z))]\\ &-\lambda_{GP}\mathbb{E}_{\hat{x}\sim p_{\hat{x}}}\left[(\|\nabla_{\hat{x}}D(\hat{x})\|_{2}-1)^{2}\right]\end{split}$ (16)
-    where $G_{\rightarrow}$ represents the cascade of generators, $D$ is the discriminator, $x$ is a sample from the real data distribution $p_{data}$, $z$ is a sample from the prior distribution $p_{z}$, $G_{\rightarrow}(z)$ is the generated sample from the cascade of generators, $\hat{x}$ is a sample from the data distribution used for the gradient penalty $p_{\hat{x}}$, $\nabla_{\hat{x}}D(\hat{x})$ is the gradient of the discriminator concerning $\hat{x}$, $\|\nabla_{\hat{x}}D(\hat{x})\|_{2}$ denotes the L2 norm of this gradient, and $\lambda_{GP}$ is the gradient penalty coefficient.
-The term $(\|\nabla_{\hat{x}}D(\hat{x})\|_{2}-1)^{2}$ enforces the Lipschitz constraint on the discriminator.
-  5. 5.
-    Auxiliary Loss Integration: Each generator $G_{i}$ minimizes a combination of the discriminator loss and its corresponding auxiliary loss, scaled by a coefficient $\lambda_{AL_{i}}$.
-  6. 6.
-    Final Output: The final synthetic dataset $\hat{X}$ is generated by passing a noise vector $z$ through the entire cascade of generators.
-IV Mutual Informatiaon
-Mutual Information (MI) measures the mutual dependence between two random variables (linear or non-linear), initially introduced by Shannon in his foundational work on information theory [94].
-It quantifies the amount of information gained about one random variable through the observation of another [95].
-Unlike the correlation coefficient, limited to real-valued random variables and linear relationships, MI is more general, capturing both linear and non-linear dependencies [96].
-The concept of MI is closely related to entropy, a fundamental notion in information theory that measures the expected amount of information in a random variable [97].
-MI can be understood as the expected value of the pointwise mutual information (PMI), reflecting how much the joint distribution of two variables deviates from the product of their marginal distributions [98].
-IV-A Comparison with Other Methods
-The rationale for selecting MI over other methods lies in its comprehensive mathematical framework.
-Initially, this study considered two general approaches for choosing the best features based on their dependence on the target feature and their relationships.
-Correlation was the first and most obvious choice, mathematically defined as the Pearson correlation coefficient (equ: LABEL:EQ:_CORR).
-However, correlation only accounts for linear relationships [99].
-In real-world scenarios, non-linear relationships often need to be identified and quantified to select the optimal set of features.
-$\text{Corr}(X,Y)=\frac{\text{Cov}(X,Y)}{\sigma_{X}\sigma_{Y}}$ (17)
-where $\text{Cov}(X,Y)$ is the covariance between variables $X$ and $Y$, and $\sigma_{X}$ and $\sigma_{Y}$ are the standard deviations of $X$ and $Y$, respectively.
-While correlation is effective for capturing linear relationships, it does not account for the non-linear relationships that often exist in real-world data, limiting its effectiveness for feature selection.
-The second option was a tree-based AI technique, known for its robustness and ability to identify essential features rigorously [100].
-However, tree-based methods suffer from the ”black box” problem, offering little interpretability regarding why certain features are selected.
-While approaches like Information Gain (IG) have been used for feature selection, as demonstrated by one study [101], Information Gain only measures the relationship between individual features and the target variable, overlooking interdependencies between features [102].
-Mathematically, it measures the reduction in entropy when a feature $X$ is used to predict the target variable $Y$.
-$IG(Y,X)=H(Y)-H(Y|X)$ (18)
-where $H(Y)$ is the entropy of the target variable $Y$, and $H(Y|X)$ is the conditional entropy of $Y$ given $X$.
-Although Information Gain quantifies the relationship between individual features and the target variable, it ignores interdependencies among the features.
-In contrast, MI captures the relationship between features and the target and the dependencies among the features themselves, providing a more holistic evaluation [103]. MI is defined as:
-$MI(X,Y)=H(X)+H(Y)-H(X,Y)$ (19)
-where $H(X)$ and $H(Y)$ are the entropies of variables $X$ and $Y$, and $H(X,Y)$ is their joint entropy.
-By capturing non-linear dependencies and interactions among features, MI offers a more comprehensive framework for feature selection.
-Compared to other AI-based feature selection methods, such as recursive feature elimination or embedded methods based on decision trees [104], MI has the upper hand by being model-agnostic.
-Many AI-based techniques are tied to specific algorithms or models, which may introduce biases or limit generalizability across different datasets or classifiers.
-MI, however, evaluates feature importance based on statistical dependencies independent of any specific model, making it a versatile and unbiased approach [105].
-This allows MI to provide a more robust selection process, especially in scenarios where feature relationships are complex and nonlinear, a common occurrence in cybersecurity network data.
-IV-B Application of Mutual Information in this Study
-In this study, MI is crucial for measuring the dataset’s dependency between different features (variables).
-It helps understand the strength of relationships between variables, which is essential for modeling and analysis.
-For example, mutual information was used to quantify the relation between each feature $x_{i}$ and the target $y$ and between each feature to identify interdependencies within the features.
-This step was crucial to filter down the essential features that explain the target variables and to cluster the features exhibiting similar behavior.
-The mutual information score of every feature was determined concerning the target, and the top 25% of the features were selected based on the mutual information weights (Table III & Table IV).
-This way, the computation was less intensive, the information was high, and the results were more interpretable.
-V Evaluation Metrics
-The evaluation metrics utilized in this study align with those used in previous studies, as outlined in section II of this paper.
-The following section details how these metrics were employed in this study and how each metric was measured to assess the experimental results.
+The generation of synthetic network traffic data has gained significant attention due to challenges associated with using real-world datasets, such as privacy concerns, ethical limitations, and data imbalance.
+Existing approaches for synthetic data generation can broadly be classified into Non-AI methods and State-of-the-Art (SOTA) AI methods.
+  • Non-AI Methods for Synthetic Data Generation:
+    Non-AI approaches primarily focus on statistical techniques and rule-based models for generating synthetic tabular data.
+Common methods include ROS (Random Over-Sampling), SMOTE (Synthetic Minority Over-sampling Technique), ADASYN (Adaptive Synthetic Sampling), GMM (Gaussian Mixture Models), and Cluster Centroids (CC) [22].
+These methods are often applied in scenarios addressing class imbalance rather than comprehensive traffic simulation [36].
+For instance, SMOTE and ADASYN have been used to oversample minority classes within network traffic datasets and are frequently compared with modern AI methods such as Variational Autoencoders (VAEs) and GANs (Generative Adversarial Networks) [37].
+    Hybrid approaches such as SMOTE-GAN, COPULA-GAN, and Bayesian GAN (GANBLR) [38, 39] have also been explored, combining statistical and generative AI techniques to enhance synthetic data fidelity [40, 41, 42].
+Despite their effectiveness in managing class imbalance, these approaches fall short in capturing the complex temporal and structural dependencies of network traffic [21, 43].
+  • State-of-the-Art AI Methods:
+    AI-based methods dominate the field of synthetic data generation, particularly in recent years. Techniques such as GANs, VAEs, and Diffusion Models have emerged as the leading approaches.
+    • GANs (Generative Adversarial Networks): GAN-based models like CTGAN [32], WGAN with gradient penalty [16, 44], CASTGAN [35], GANBLR++ [38, 34] and have been widely used for synthetic tabular data generation.
+For network traffic-specific datasets, models such as FlowGAN [4], STAN [3], NetShare [6], DoppelGANger [45], and 5GT-GAN [46], MirageNet [5], have shown significant promise.
+    • VAEs (Variational Autoencoders): VAEs, particularly TVAE (Tabular VAE), are used as alternatives to GANs for tabular data generation.
+Recent studies, such as XIDINTFL-VAE, have demonstrated the application of VAEs in handling class imbalance within network traffic datasets [37].
+    • Diffusion Models: These models, including TABDDPM [31], NetDiffus [47], and NetDiffusion [7], represent an emerging paradigm in synthetic data generation.
+The recent NetDiffus framework has outperformed GAN-based methods, providing a 66.4% increase in the fidelity of generated data and an 18.1% increase in downstream ML tasks.
+They have been reported to outperform GANs and VAEs in terms of data fidelity and diversity.
+    • Simulators and Transformer Models: A recent study examined the feasibility of state space models, specifically Mamba, for generating packet-level synthetic network traffic data [48].
+They compared the simulator with the state-of-the-art AI models, including a transformer-based TrafficGPT [49].
+While these methods have individually shown great potential, they often lack direct comparative analysis under standardized evaluation metrics, making it difficult to identify the most suitable approach for specific network traffic data and its applications.
+Thus, a significant limitation in the current body of research is the absence of direct comparative studies evaluating Non-AI and AI (conventional and generative/SOTA) methods under unified experimental settings and standardized evaluation scenarios.
+From the reviewed literature, the following key research gaps are identified:
+  • Limited comparative analysis of non-AI and AI (conventional and generative/SOTA) methods.
+  • Lack of standard evaluation metrics to assess synthetic data quality across different approaches.
+  • Class imbalance issues remain unexplored in the context of network traffic datasets.
+This study aims to address these gaps by systematically comparing non-AI and AI (conventional and generative/SOTA) methods using standardized evaluation metrics of fidelity and utility, with a specific focus on handling class imbalance in synthetic network traffic datasets.
+Therefore, for the comprehensive comparative study, we have selected twelve prominent synthetic tabular data methods.
+Technical details of these methods are in the next section.
+III Synthetic Network Traffic Data Generation Methods
+The generation of synthetic network traffic data can be broadly categorized into Non-AI (statistical) methods and AI-based methods, which include both classical AI and generative AI approaches (Figure 1).
+This section provides a structured overview of these methods and explains the comparative evaluation approach used in this study.
+III-A Overview of Synthetic Data Generation Methods
+Non-AI Methods (Statistical Approaches): These methods rely on mathematical models to generate synthetic data by manipulating existing samples or approximating underlying distributions.
+They are widely used due to their simplicity and low computational cost but often struggle to capture complex traffic dependencies [22, 50].
+AI-Based Methods (Classical AI and Generative AI): AI-driven approaches aim to model network traffic more effectively by learning patterns from real data.
+Classical AI methods use probabilistic models and neural architectures to approximate traffic distributions, while generative AI techniques, such as GANs and diffusion models, focus on generating highly realistic synthetic data through adversarial training or iterative refinements [51, 52].
+A summary of these methods, including their key features, strengths, and limitations, is presented in Table I.
+These methods were selected based on their prevalence in prior research and their relevance to the data structure of network traffic.
+This study focuses on comparing these approaches to determine which method best preserves fidelity, maintains class balance, and remains computationally efficient in generating synthetic network traffic data.
+III-B Comparative Evaluation Approach
+Rather than solely describing these methods, we systematically compare them based on the following key evaluation criteria:
+  • Fidelity: The degree to which synthetic data retains real network traffic’s statistical and temporal characteristics.
+  • Utility: The effectiveness of synthetic data in training machine learning models for network traffic classification.
+  • Class Balance: The ability of each method to generate data that reflects the true distribution of network events, addressing imbalanced datasets.
+  • Scalability: The computational efficiency of each method and its feasibility for large-scale synthetic data generation.
+Each method is applied to the NSL-KDD and CIC-IDS2017 datasets to ensure an objective and standardized comparison [53, 22].
+Performance is evaluated using quantitative metrics aligned with these criteria, and the results are analyzed in later sections.
+This structured comparison allows us to determine the trade-offs between different approaches and provide practical recommendations for selecting the most suitable synthetic data generation techniques based on specific application needs.
+IV Evaluation Metrics
+The following section details how each evaluation metric was measured to assess the experimental results.
+Figure 2: Correlation Heatmap for NSL KDD (Part 1/2)
+Figure 3: Correlation Heatmap for NSL KDD (Part 2/2)
   1. 1.
@@
 Failing to maintain these structural properties can lead to synthetic data that, while statistically similar on a superficial level, fails to capture the logical and contextual nuances necessary for accurate downstream analyses and applications [57].
-In this study, the following metrics were used to measure the statistical similarity (fidelity), as shown in Tables V to X:
-    1. (a)
-      Data Boundaries (DB): This metric checks whether the synthetic data adheres to the original data’s logical minimum and maximum values.
-For example, a binary column should contain only binary values, ensuring no out-of-bound entries.
-    2. (b)
-      Correlation (Corr): This is assessed by examining the correlation heatmap of all variables’ absolute values of the correlation coefficients.
-It evaluates the strength and direction of relationships in both real and synthetic datasets.
-    3. (c)
-      Probability Distribution (PD): This compares the underlying probability distribution of each variable between the real and synthetic datasets, ensuring that their statistical properties are aligned.
-  2. 2.
-    Performance (utility/accuracy) is a widely used metric to evaluate the utility of synthetic data, particularly in the context of ML and DL models [109].
-This metric compares how models trained on synthetic data perform relative to those trained on real data [110].
-The most common approach is to contrast TRTR (train and test on real data) with TSTR (train on synthetic data and test on real data).
-By assessing the model’s accuracy, F1 score, precision, and recall, researchers can determine the efficacy of the synthetic data.
-Suppose the model’s performance metrics in the TSTR scenario are comparable to or exceed those in the TRTR scenario.
-In that case, the synthetic data is considered a viable alternative to the real data [14].
-    Moreover, a well-performing TSTR indicates that the synthetic data captures the essential patterns and relationships within the real data, making it suitable for model development, validation, and deployment in privacy-sensitive environments.
-However, it’s crucial to note that synthetic data’s utility is not solely dependent on performance metrics; it also relies on the data’s ability to generalize across different models and tasks, which should be considered in comprehensive evaluations [111].
-    Thus, in this study, the accuracy of TRTR was compared with TSTR for measuring ML utility, as reported in Tables V to X.
-  3. 3.
-    Class Balance (CB) is a critical metric for evaluating the quality of synthetic tabular data, particularly in classification tasks.
-Class balance refers to the distribution of instances across different classes in the dataset[112].
-Maintaining a balanced class distribution in real-world datasets is essential for training robust ML models, as imbalanced courses can lead to biased models that perform poorly on underrepresented classes [113].
-    Ensuring that the class distribution closely mirrors the real data is essential when assessing synthetic data.
-This involves not only comparing the overall proportions of each class but also examining the stability of these proportions across different subsets of the data.
-Some methods and tools for generating synthetic data inherently include mechanisms to evaluate and maintain class balance, ensuring that the synthetic dataset accurately reflects the class distribution of the real dataset.
-Suppose a synthetic data generation method does not inherently address class balance.
-In that case, it may require additional post-processing to adjust class distributions and avoid introducing bias into the model training process [114].
-    Thus, the CB results mentioned in Tables V to X represent how much difference exists between the NORMAL and ATTACK cases.
-If the classes in the synthetic data are the same, there is zero difference; otherwise, the difference is detailed in the results table later in this paper.
-VI Experiment Setup
-TABLE I: Summary of 26 Encoded selected Features from NSL-KDD Dataset
+In this study, the following metrics were used to measure the statistical similarity (fidelity), as shown in results Table VI and Table VII:
+    Figure 4: Probability Distribution Comparison for NSL KDD.
+    IV-A Experiment Setup
+    TABLE II: Summary of 26 Encoded selected Features from NSL-KDD Dataset [58]
     S.No Feature Description
@@
     26 target Class label indicating if the connection is normal or an attack
-TABLE II: Summary of 21 selected Features from CICIDS-17 Dataset
+    TABLE III: Summary of 21 selected Features from CIC-IDS2017 Dataset [12]
     S.No Feature Description
@@
     21 target Class label indicating if the flow is benign or malicious
-TABLE III: Mutual Information Score of NSL-KDD Features
+    1. (a)
+      Data Structure (DS): This metric checks whether the synthetic data adheres to the original data’s logical minimum and maximum values.
+For example, a binary column should contain only binary values, ensuring no out-of-bound entries.
+      Table VI and Table VII include a column where each value is calculated based on whether all binary and boolean variables adhere to their expected values.
+This means that binary variables should be either 0 or 1, and boolean variables should be either TRUE or FALSE.
+If the synthetic data generated by a method does not meet these structural requirements, the result will be marked as NO.
+      For example, in Table VI, the first synthetic data set was generated by ROS, and the value in the DS column is YES, indicating that all binary and boolean variables adhered to their expected values.
+However, the next row, corresponding to SMOTE, shows NO under the DS column because, apart from the target variable in the NSL-KDD data, the binary and boolean variables did not conform to their expected values.
+This evaluation method helps assess whether each approach maintains the required data boundaries.
+    2. (b)
+      Correlation (Corr): This assessment is performed by examining the correlation heatmap, which represents the absolute values of the correlation coefficients for all variables (see Figure 2 and Figure 3 for key NSL KDD heatmaps).
+It evaluates the strength and direction of relationships in both real and synthetic datasets.
+If there is a significant difference between the absolute correlation heatmap of the synthetic data and the real data, as well as in the heatmap of the absolute correlation differences, the result is marked as NO.
+Otherwise, it is marked as YES in the Corr column of all of the results tables.
+    3. (c)
+      Probability Distribution (PD): This compares the underlying probability distribution of each variable between the real and synthetic datasets, ensuring that their statistical properties are aligned.
+      To calculate the statistics for the PD comparison, the following steps were undertaken:
+      1. i.
+        For every variable in the synthetic and real datasets, the underlying PD was estimated using Kernel Density Estimation (KDE).
+      2. ii.
+        Each variable’s PD was visualized.
+For clarity, individual figures showing the real and synthetic PDs were generated.
+In some cases, side-by-side figures were created to provide a direct comparison (Figure 2).
+Complete PD comparison graphs for both NSL-KDD and CIC-IDS2017 are provided in Appendix A.
+      3. iii.
+        The alignment of the PD was quantified using the following formula:
+[…]
+        where “Number of Variables with Different PD indicates the count of variables whose PD differed between the real and synthetic datasets.
+      For the NSL-KDD dataset, which contains 26 total variables, the results are summarized in Table VI.
+Under the PD column, for the ROS method, all variables exhibited identical PDs according to distribution curves in Figure 4, resulting in a PD of $0\%$ difference.
+Conversely, for the Adaptive Synthetic Sampling method, one variable out of 26 exhibited a different PD (see Appendix A for a complete graphical PD comparison).
+Using Equation 1, the percentage difference was calculated as:
+[…]
+  2. 2.
+    Performance (utility/accuracy) is a widely used metric to evaluate the utility of synthetic data, particularly in the context of ML and DL models [59].
+This metric compares how models trained on synthetic data perform relative to those trained on real data [60].
+The most common approach is to contrast TRTR (train and test on real data) with TSTR (train on synthetic data and test on real data).
+By assessing the model’s accuracy, F1 score, precision, and recall, researchers can determine the efficacy of the synthetic data.
+Suppose the model’s performance metrics in the TSTR scenario are comparable to or exceed those in the TRTR scenario.
+In that case, the synthetic data is considered a viable alternative to the real data [38].
+    Moreover, a well-performing TSTR indicates that the synthetic data captures the essential patterns and relationships within the real data, making it suitable for model development, validation, and deployment in privacy-sensitive environments.
+However, it’s crucial to note that synthetic data’s utility is not solely dependent on performance metrics; it also relies on the data’s ability to generalize across different models and tasks, which should be considered in comprehensive evaluations [61].
+    Thus, in this study, the accuracy of TRTR was compared with TSTR for measuring ML utility, as reported in Table VI and Table VII.
+  3. 3.
+    Class Balance (CB) is a critical metric for evaluating the quality of synthetic tabular data, particularly in classification tasks.
+Class balance refers to the distribution of instances across different classes in the dataset[62].
+Maintaining a balanced class distribution in real-world datasets is essential for training robust ML models, as imbalanced courses can lead to biased models that perform poorly on underrepresented classes [63].
+    Class Balance Difference (%)=|$\begin{split}\text{Class Balance Difference (\%)}=&\left|P_{\text{synthetic},\text{Normal}}-P_{\text{synthetic},\text{Attack}}\right|\\ &\times 100\end{split}$ (3)
+    Thus, the CB results mentioned in Table VI and Table VII represent how much difference exists between the NORMAL and ATTACK cases.
+If the classes in the synthetic data are the same, there is zero difference; otherwise, the difference percentage is calculated based on the equation.
+This equation 3 gives 0% diff if the normaland attackclasses are equal, i.e., follow a 50-50 ratio in the whole data.
+Thus, in Table VI, under the CB column for ROS and SMOTE, it is mentioned 0% diff, and for the Adaptive Synthetic Sampling, it is 0.14% diff.
+V Experimental Setup
+V-A Mutual Informatiaon
+Mutual Information (MI) measures the mutual dependence between two random variables (linear or non-linear), initially introduced by Shannon in his foundational work on information theory [64].
+It quantifies the amount of information gained about one random variable through the observation of another [65].
+Unlike the correlation coefficient, limited to real-valued random variables and linear relationships, MI is more general, capturing both linear and non-linear dependencies [66].
+The concept of MI is closely related to entropy, a fundamental notion in information theory that measures the expected amount of information in a random variable [67].
+MI can be understood as the expected value of the pointwise mutual information (PMI), reflecting how much the joint distribution of two variables deviates from the product of their marginal distributions [68].
+V-A1 Comparison with Other Methods
+The rationale for selecting MI over other methods lies in its comprehensive mathematical framework.
+Initially, this study considered two general approaches for choosing the best features based on their dependence on the target feature and their relationships.
+Correlation was the first and most obvious choice, mathematically defined as the Pearson correlation coefficient (equ: 4).
+However, correlation only accounts for linear relationships [69].
+In real-world scenarios, non-linear relationships often need to be identified and quantified to select the optimal set of features.
+$\text{Corr}(X,Y)=\frac{\text{Cov}(X,Y)}{\sigma_{X}\sigma_{Y}}$ (4)
+where $\text{Cov}(X,Y)$ is the covariance between variables $X$ and $Y$, and $\sigma_{X}$ and $\sigma_{Y}$ are the standard deviations of $X$ and $Y$, respectively.
+While correlation is effective for capturing linear relationships, it does not account for the non-linear relationships that often exist in real-world data, limiting its effectiveness for feature selection.
+The second option was a tree-based AI technique, known for its robustness and ability to identify essential features rigorously [70].
+However, tree-based methods suffer from the ”black box” problem, offering little interpretability regarding why certain features are selected.
+While approaches like Information Gain (IG) have been used for feature selection, as demonstrated by one study [71], Information Gain only measures the relationship between individual features and the target variable, overlooking interdependencies between features [72].
+Mathematically, it measures the reduction in entropy when a feature $X$ is used to predict the target variable $Y$.
+$IG(Y,X)=H(Y)-H(Y|X)$ (5)
+where $H(Y)$ is the entropy of the target variable $Y$, and $H(Y|X)$ is the conditional entropy of $Y$ given $X$.
+Although Information Gain quantifies the relationship between individual features and the target variable, it ignores interdependencies among the features.
+In contrast, MI captures the relationship between features and the target and the dependencies among the features themselves, providing a more holistic evaluation [73]. MI is defined as:
+$MI(X,Y)=H(X)+H(Y)-H(X,Y)$ (6)
+where $H(X)$ and $H(Y)$ are the entropies of variables $X$ and $Y$, and $H(X,Y)$ is their joint entropy.
+By capturing non-linear dependencies and interactions among features, MI offers a more comprehensive framework for feature selection.
+Compared to other AI-based feature selection methods, such as recursive feature elimination or embedded methods based on decision trees [74], MI has the upper hand by being model-agnostic.
+Many AI-based techniques are tied to specific algorithms or models, which may introduce biases or limit generalizability across different datasets or classifiers.
+MI, however, evaluates feature importance based on statistical dependencies independent of any specific model, making it a versatile and unbiased approach [75].
+This allows MI to provide a more robust selection process, especially in scenarios where feature relationships are complex and nonlinear, a common occurrence in network data.
+V-A2 Application of Mutual Information in this Study
+In this study, MI is crucial for measuring the dataset’s dependency between different features (variables).
+It helps understand the strength of relationships between variables, which is essential for modeling and analysis.
+For example, mutual information was used to quantify the relation between each feature $x_{i}$ and the target $y$ and between each feature to identify interdependencies within the features.
+This step was crucial to filter down the essential features that explain the target variables and to cluster the features exhibiting similar behavior.
+The mutual information score of every feature was determined concerning the target, and the top 25% of the features were selected based on the mutual information weights (Table IV & Table V).
+This way, the computation was less intensive, the information was high, and the results were more interpretable.
+V-B Datasets & Pre-processing
+  1. 1.
+    NSL-KDD dataset, a refined version of the original KDD Cup 1999 dataset, is widely used in cybersecurity to evaluate IDS [58].
+It addresses some critical issues present in its predecessor, such as redundant records and imbalanced distribution, thereby providing a more accurate and reliable benchmark for performance evaluation.
+For this study, the training portion of the NSL-KDD dataset, originally consisting of 125,973 instances and 42 columns (41 features and 1 (binary) target), was used.
+    The preprocessing of the NSL-KDD dataset involved several essential steps to prepare it for synthetic data generation and evaluation.
+Initially, categorical columns were encoded to transform non-numeric data into a suitable format for ML algorithms.
+This encoding process ensured that all categorical variables were converted into numerical values, facilitating the subsequent analysis.
+Following the encoding, feature extraction was performed to identify the most significant features using Mutual Information.
+This process selected the features’ top 25% (Q1) based on the mutual information weights.
+This step resulted in a reduced feature set of 26 encoded features (Table II), which were the most important (based on the information weights; Table IV) for accurately representing the data.
+This refined feature set was then used in the data generation phase, ensuring that the synthetic data closely mirrored the critical characteristics of the original dataset.
+  2. 2.
+    CIC-IDS2017 dataset, part of the Canadian Institute for Cybersecurity’s intrusion detection dataset collection, is widely employed for evaluating cybersecurity systems [12].
+It provides a comprehensive network traffic data set with diverse attack types and normal traffic, offering a robust benchmark for performance assessment.
+CIC-IDS2017 has one week of captured network traffic data in eight comma-separated files (CSV).
+These files are generated based on the attack types.
+There are 2,830,743 instances of all eight files and 79 columns (78 numerical features and one binary target).
+    Preprocessing of the CIC-IDS2017 dataset involved several vital steps to ready it for synthetic data generation and evaluation.
+First, all the files were individually tested for null values and duplicated columns.
+In this process, it was found that there were some infinity values in the data.
+Thus, all the null values, infinity, and duplicated column ”Fwd Header Length” were removed.
+Secondly, all the cleaned data files were concatenated into a single CSV file for further feature selection.
+Lastly, mutual information was performed to select the variables’ top 25% (Q1) based on the mutual information weights.
+This step resulted in a reduced feature set of 21 features (Table III), which were determined to be the most important (based on the information weights; V) for accurately representing the data.
+This reduced feature set was then employed in the synthetic data generation phase, ensuring that the synthetic data accurately reflected the essential characteristics of the original dataset.
+TABLE IV: Mutual Information Score of NSL-KDD Features
 Feature Mutual Information Score
@@
 service_domain_u 0.048430
-TABLE IV: Mutual Information Score of CICIDS-17 Features
+TABLE V: Mutual Information Score of CIC-IDS2017 Features
 Feature Mutual Information Score
@@
 Flow Bytes/s 0.209784
-VI-A Hardware Setup
+V-B1 Hardware Setup
 The experiments were performed on a workstation with a 13th Gen Intel® Core™ i9-13900 processor.
@@
 A dedicated GPU, NVIDIA GeForce RTX 4090, was used for these experiments.
-VI-B Software Setup
+V-B2 Software Setup
 The software environment for the experiments was based on Microsoft Windows 11 Pro.
@@
 PGMPy and SciPy were also used for probabilistic graphical models and scientific computations.
-VI-C Datasets & Pre-processing
-  1. 1.
-    NSL-KDD dataset, a refined version of the original KDD Cup 1999 dataset, is widely used in cybersecurity to evaluate IDS [115].
-It addresses some critical issues present in its predecessor, such as redundant records and imbalanced distribution, thereby providing a more accurate and reliable benchmark for performance evaluation.
-For this study, the training portion of the NSL-KDD dataset, originally consisting of 125,973 instances and 42 columns (41 features and 1 (binary) target), was used.
-    The preprocessing of the NSL-KDD dataset involved several essential steps to prepare it for synthetic data generation and evaluation.
-Initially, categorical columns were encoded to transform non-numeric data into a suitable format for ML algorithms.
-This encoding process ensured that all categorical variables were converted into numerical values, facilitating the subsequent analysis.
-Following the encoding, feature extraction was performed to identify the most significant features using Mutual Information.
-This process selected the features’ top 25% (Q1) based on the mutual information weights.
-This step resulted in a reduced feature set of 26 encoded features (Table I), which were the most important (based on the information weights; Table III) for accurately representing the data.
-This refined feature set was then used in the data generation phase, ensuring that the synthetic data closely mirrored the critical characteristics of the original dataset.
-  2. 2.
-    CICIDS-17 dataset, part of the Canadian Institute for Cybersecurity’s intrusion detection dataset collection, is widely employed for evaluating cybersecurity systems [18].
-It provides a comprehensive network traffic data set with diverse attack types and normal traffic, offering a robust benchmark for performance assessment.
-CICIDS-17 has one week of captured network traffic data in eight comma-separated files (CSV).
-These files are generated based on the attack types.
-There are 2,830,743 instances of all eight files and 79 columns (78 numerical features and one binary target).
-    Preprocessing of the CICIDS-17 dataset involved several vital steps to ready it for synthetic data generation and evaluation.
-First, all the files were individually tested for null values and duplicated columns.
-In this process, it was found that there were some infinity values in the data.
-Thus, all the null values, infinity, and duplicated column ”Fwd Header Length” were removed.
-Secondly, all the cleaned data files were concatenated into a single CSV file for further feature selection.
-Lastly, mutual information was performed to select the variables’ top 25% (Q1) based on the mutual information weights.
-This step resulted in a reduced feature set of 21 features (Table II), which were determined to be the most important (based on the information weights; IV) for accurately representing the data.
-This reduced feature set was then employed in the synthetic data generation phase, ensuring that the synthetic data accurately reflected the essential characteristics of the original dataset.
-VII Experimental Results and Discussion
-TABLE V: Comparison of Synthetic Data by Standard Methods on NSL-KDD Dataset
-Method Description Pros Cons Performance Summary
-SS CB Accuracy
+VI Results and Discussion
+TABLE VI: Comparison of Synthetic Data on NSL-KDD Dataset
+Category Method SS CB Accuracy
 DS Corr PD
-ROS [116] Randomly duplicates minority class samples Simple, low cost Overfitting risk, no new samples Yes Yes 0% diff 0% diff 0.9998
-SMOTE [63] Generates synthetic samples by interpolation Reduces overfitting, better minority representation May create ambiguous samples, high computation No Yes 0% diff 0% diff 0.9999
-Adaptive Synthetic Sampling [67] Adjusts sampling based on data distribution Flexible, effective for imbalance Complex tuning, high computation No Yes 3.8% diff 0.14% diff 0.9998
-Cluster based oversampling [117] Oversamples within identified clusters Captures local structures Sensitive to clustering algorithm No Yes 0% diff 0% diff 0.9995
-Gaussian Mixture Model [75] Uses Gaussian distributions for sampling Captures complex distributions Assumes data distribution No Yes 99.1% diff 99.9% diff 0.5314
-SS” = Statistical Similarity, ”DS” = Data Structure, ”Corr” = Correlation, ”PD” = Probability Distribution, ”CB” = Class Balance, ”Accuracy” = Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data = 0.9995, ”TSTR” = Train on Synthetic and Test on Real Data
-TABLE VI: Comparison of Synthetic Data by Standard Methods on CICIDS-17 Dataset
-Method Statistical Similarity Class Balance Accuracy (0.9978)
-Data Structure Correlation Probability Distribution
-ROS Yes Yes 0% diff 0% diff 0.9991
-SMOTE Yes Yes 0% diff 0% diff 0.9990
-Adaptive Synthetic Sampling Yes Yes 0% diff 0.2% diff 0.9986
-Cluster-based oversampling No Yes 0% diff 0% diff 0.9930
-Gaussian Mixture Model No No 61.9% diff 46.1% diff 0.4509
-Accuracy” = Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
-TABLE VII: Comparison of Synthetic Data by AI Methods on NSL-KDD Dataset
-Method Description Pros Cons Performance Summary
-SS CB Accuracy
-DS Corr PD
-Bayesian Neural Networks (BN) [118] Uses Bayesian inference for uncertainty estimation Quantifies uncertainty, probabilistic dependencies Requires prior knowledge, complex tuning No No 0% diff 17.2% diff 0.9587
-Variational Tabular Autoencoders (TVAEs)[119] Neural networks for unsupervised feature learning Learns representations, reduces dimensionality KL divergence issue, risk of overfitting Yes Yes 23.1% diff 26.7% diff 0.9807
-Diffusion Models [86] Uses diffusion for data generation Captures temporal dependencies Needs sequential data, high computational cost No No 98.3% diff 34.2% diff 0.51
-Conditional Tabular Generative Adversarial Networks (CTGAN) [12] Adversarial training between generator and discriminator High-quality data, models complex distributions Training instability, mode collapse Yes Yes 11.5% diff 5.7% diff 0.9640
-SS = Statistical Similarity, DS = Data Structure, Corr = Correlation, PD = Probability Distribution, CB = Class Balance, ”Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data = 0.9995, ”TSTR” = Train on Synthetic and Test on Real Data
-TABLE VIII: Comparison of Synthetic Data by AI methods on CICIDS-17 Dataset
-Method Statistical Similarity Class Balance Accuracy (0.9978)
-Data Structure Correlation Probability Distribution
-Bayesian Neural Networks (BN)
-Tabular Variational Autoencoder (TVAE) Yes No 1.4% diff 69.4% diff 0.9718
-Conditional Tabular GAN (CTGAN) Yes Yes 0.8% diff 5.2% diff 0.9603
-Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
-TABLE IX: Comparison of Synthetic Data by prominent GANs on NSL-KDD Dataset
-Method Statistical Similarity Class Balance Accuracy (0.9995)
-Data Structure Correlation Probability Distribution
+Non-AI (Statistical) ROS Yes Yes 0% diff 0% diff 0.9998
+SMOTE No Yes 0% diff 0% diff 0.9999
+ADASYN No Yes 3.8% diff 0.14% diff 0.9998
+CC No Yes 0% diff 0% diff 0.9995
+GMM No Yes 99.1% diff 99.9% diff 0.5314
+AI-Based (Classical + Generative) BN No No 0% diff 17.2% diff 0.9587
+TVAE Yes Yes 23.1% diff 26.7% diff 0.9807
+TABDDPM No No 98.3% diff 34.2% diff 0.51
 CTGAN Yes Yes 7.7% diff 6.7% diff 0.9667
@@
 CasTGAN No Yes 27% diff 11.1% diff 0.9582
-Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
-TABLE X: Comparison of Synthetic Data by prominent GANs on CICIDS-17 Dataset
-Method Statistical Similarity Class Balance Accuracy (0.9978)
-Data Structure Correlation Probability Distribution
+SS” = Statistical Similarity, ”DS” = Data Structure, ”Corr” = Correlation, ”PD” = Probability Distribution, ”CB” = Class Balance, ”Accuracy” = Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data = 0.9995, ”TSTR” = Train on Synthetic and Test on Real Data
+TABLE VII: Comparison of Synthetic Data on CIC-IDS2017 Dataset
+Category Method SS CB Accuracy
+DS Corr PD
+Non-AI (Statistical) ROS Yes Yes 0% diff 0% diff 0.9991
+SMOTE Yes Yes 0% diff 0% diff 0.9990
+ADASYN Yes Yes 0% diff 0.2% diff 0.9986
+CC No Yes 0% diff 0% diff 0.9930
+GMM No No 61.9% diff 46.1% diff 0.4509
+AI-Based (Classical + Generative) BN
+TVAE Yes No 1.4% diff 69.4% diff 0.9718
+TABDDPM
 CTGAN Yes Yes 0% diff 5% diff 0.95381
@@
 CasTGAN No No 4.8% diff 58.3% diff 0.9734
-Accuracy” Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data, ”TSTR” = Train on Synthetic and Test on Real Data
-Given the surge in interest in generating synthetic tabular data, the GAN framework has emerged as a seminal approach, attracting substantial research attention.
-Consequently, this paper delves deeply into evaluating current high-performing GAN models, leveraging recent advancements in this domain.
-A meticulous analysis of four leading models is presented in Section III, elucidating their methodologies and performance metrics.
-VII-A Results
-This study evaluated various methods for generating synthetic tabular data, covering both statistical and AI-based approaches.
-VII-A1 Statistical Methods
-Five widely recognized statistical methods were assessed, as outlined in the survey by [3] (detailed in Section III). The findings indicate the following:
-  • ROS emerged as the most effective statistical method for preserving the data distribution and ensuring class balance. It maintained a high degree of statistical similarity to the original dataset.
-  • SMOTE and Cluster Centroid outperformed ROS by maintaining statistical similarity and enhancing ML utility, achieving a $99\%$ accuracy while preserving class balance.
-  • GMM showed strong performance in capturing feature correlations but underperformed in statistical similarity and ML utility, with an accuracy of only $53\%$.
-  • The results were consistent across both the synthetic NSL-KDD and CICIDS-17 datasets (Table V and Table VI).
-VII-A2 AI Methods
-Four prominent AI and generative AI methodologies were evaluated (details in Section III). The results are summarized as follows:
-  • Bayesian Networks struggled with maintaining correlation and structure in the NSL-KDD dataset, especially for binary features. Diffusion models had the lowest ML utility accuracy at $51\%$.
-  • TVAE and Conditional Tabular CTGAN performed exceptionally well, preserving statistical properties and achieving high ML utility accuracy rates of $98\%$ and $96\%$, respectively (Table VII).
-  • Consistent performance was observed with the CICIDS-17 dataset, where TVAE slightly outperformed CTGAN in ML utility.
-VII-A3 Prominent GANs
-Four leading GAN models were evaluated in depth (detailed in Section III). The results showed:
-  • CTGAN and Copula-GAN outperformed other models, with ML utility accuracy scores of $97\%$ and $98\%$, respectively, on the NSL-KDD and CICIDS-17 datasets.
-  • CastGAN showed strong performance in handling feature dependencies but struggled to preserve boundary conditions for specific binary columns.
-  • GANBLR++, based on Bayes’ network principles, exhibited performance metrics closely aligned with those of individual BN models.
-VII-B Discussion
-The results from this study provide several key insights into the effectiveness of various synthetic data generation methods:
-VII-B1 Statistical Methods
-While effective at maintaining data distribution, ROS is limited in generating genuinely synthetic data, making it less suitable for cybersecurity applications.
-By definition, ROS does not generate novel data points; it is just random sampling.
-SMOTE and Cluster Centroid, by enhancing ML utility, present better alternatives for applications where both statistical fidelity and predictive accuracy are critical.
-Both SMOTE and cluster centroids create novel data points by interpolation and average representation, respectively [63, 120].
-GMM’s performance highlights its strength in capturing feature correlations and creating novel data points using the fundamental definition of samples from fitted distributions.
-Still, its lower utility score suggests it may not be ideal for all applications [121, 122].
-VII-B2 AI Methods
-Bayesian Networks’ struggle with larger datasets, particularly with multi-class features, indicates a need for better handling such data structures in AI models.
-The potential reason for this struggle comes from the probabilistic graphical model.
-This graph tries to represent dependency between variables, and in the case of the CIC-IDS large dataset, it fails to create discretization because of low representation learning (memory) [123, 124].
-Diffusion models require further refinement to improve their statistical similarity and utility because they are, in general, designed for high-dimensional data with spatial relationships (e.g., images), making it challenging to capture the non-spatial, relational structure of tabular data [85].
-The success of TVAE and CTGAN emphasizes the growing importance of generative AI models in synthetic data generation.
-However, their lack of inherent class imbalance solutions suggests that additional techniques may be required to address this issue comprehensively.
-TVAE and CTGAN excel in generating high-quality synthetic tabular data due to their specialized architectures and loss functions.
-TVAE uses a probabilistic latent space to model complex dependencies, handling continuous and categorical features with different probability distributions [125].
-Its KL divergence-based loss regularizes the model, ensuring diverse yet realistic data generation [126].
-CTGAN, on the other hand, uses a conditional generator to capture feature relationships and a specific loss function that optimizes for mode-specific normalization, preserving the distributions of both majority and minority categories [127].
-The combination of these tailored architectures and well-designed loss functions enables both models to generate synthetic data that closely mirrors the statistical properties of the original dataset, making them highly effective for tabular data generation [12].
-VII-B3 Prominent GANs
-The superior performance of CTGAN and Copula-GAN reinforces their suitability for generating synthetic tabular data that retains statistical integrity and practical utility.
-CopulaGAN excels in generating synthetic tabular data because it uses copulas to model the dependencies between variables and their marginal distributions separately [128].
-This allows it to capture complex relationships between features more accurately than standard GANs.
-By decoupling the modeling of individual feature distributions from their joint dependency structure, CopulaGAN ensures that the synthetic data maintains the statistical properties of the original dataset, making it highly effective for tabular data [129].
-CastGAN’s innovative architecture of each generator and auxiliary learner for each variable shows promise, though its limitations with binary features suggest areas for future improvement [92].
-The effectiveness of GANBLR++ demonstrates the value of integrating Bayesian principles into GAN frameworks, providing a robust foundation for capturing complex probabilistic relationships in the data for both numerical and categorical datasets [90].
-VIII Future Work
+SS” = Statistical Similarity, ”DS” = Data Structure, ”Corr” = Correlation, ”PD” = Probability Distribution, ”CB” = Class Balance, ”Accuracy” = Machine Learning Utility on TRTR vs TSTR, ”TRTR” = Train and Test on Real Data = 0.9995, ”TSTR” = Train on Synthetic and Test on Real Data
+This section presents an in-depth comparative evaluation of twelve synthetic data generation methods applied to the NSL-KDD and CIC-IDS2017 datasets.
+The analysis is structured around four key objectives: fidelity, utility, class balance, and computational efficiency.
+Tables VI and VII summarize these results for both datasets.
+The evaluation focuses on two distinct categories of methods: Non-AI (Statistical) and AI-based (Classical + Generative AI).
+The performance of these methods is assessed based on statistical similarity (SS), class balance (CB), and machine learning utility (measured as accuracy in TRTR vs.
+TSTR).
+Additionally, AI-based methods’ computational cost and training efficiency are critically analyzed.
+VI-A Statistical Methods
+Statistical methods such as ROS, SMOTE, ADASYN, and Cluster Centroids (CC) are widely used for addressing class imbalance in network traffic datasets.
+However, these methods primarily rely on resampling strategies rather than truly modeling the underlying data distribution.
+While SMOTE and ADASYN generate synthetic samples, they do so through interpolation, which may fail to capture complex feature dependencies.
+Additionally, statistical oversampling techniques can lead to overfitting in downstream machine learning tasks, as the generated samples lack inherent variability compared to real network traffic.
+Prior studies [X, Y] indicate that such methods are effective in reducing bias in classification but perform poorly when high-dimensional feature interactions are critical for network anomaly detection.
+They exhibit excellent class balance maintenance, achieving 0% CB difference in most cases.
+However, these methods primarily resample existing data points rather than generating truly novel synthetic data, which limits their effectiveness in capturing complex relationships between features.
+  • SMOTE and CC outperform ROS by enhancing ML utility while preserving statistical similarity.
+  • GMM performs poorly due to its parametric limitations, struggling to model high-dimensional network traffic distributions (99.1% PD diff in NSL-KDD, 61.9% in CIC-IDS2017).
+  • Despite their fidelity limitations, statistical methods achieve near-perfect accuracy (0.999) in TRTR and TSTR settings, showing they can still be useful for ML training in balanced data scenarios.
+VI-B AI-Based Methods
+AI-based methods exhibit stronger generative capabilities but face trade-offs in fidelity, utility, and computational cost:
+  • CTGAN and CopulaGAN achieve the best balance, with high fidelity (Corr: Yes, PD diff: 7.7% (NSL-KDD), Corr: Yes, PD diff: 0% (CIC-IDS2017)) and strong utility (accuracy: 0.9667 (NSL-KDD) and 0.9538 (CIC-IDS2017)).
+  • TVAE is overall highly effective in NSL-KDD but struggles with class balance in CIC-IDS2017 (CB diff: 69.4%), indicating possible limitations in handling imbalanced network traffic distributions.
+  • Diffusion models (TABDDPM) perform poorly, particularly in high-dimensional tabular data, due to their reliance on iterative denoising rather than structured feature learning (PD diff: 98.3% (NSL-KDD)).
+VI-C Computational Cost and Training Efficiency
+The computational efficiency of AI-based methods varies significantly, impacting their scalability for large-scale network traffic simulation:
+  • Bayesian Networks (BN) failed on CIC-IDS2017 (Table VII) due to excessive memory requirements when constructing initial feature relationships.
+This highlights the challenge of applying probabilistic graphical models to large network traffic datasets.
+  • TABDDPM could not generate results for CIC-IDS2017 (Table VII), indicating prohibitively high computational costs.
+Diffusion-based models, while powerful for image data, struggle with tabular datasets due to their need for extensive forward and reverse sampling processes.
+  • CTGAN and CopulaGAN balance computational efficiency and fidelity, making them the most practical for synthetic network traffic data where scalability is critical.
+Key technical inference from the abovementioned results is that GAN-based models, such as CTGAN and CopulaGAN, perform exceptionally well because they utilize mode-specific normalization and conditional generation, which helps preserve the feature dependencies crucial for generating realistic network traffic samples.
+This ability to maintain the relationships between features contributes to their effectiveness in generating high-quality synthetic data [32, 76].
+On the other hand, diffusion models face significant challenges, particularly when applied to large tabular datasets.
+These models rely on iterative denoising, a process that becomes computationally expensive as the dataset size increases.
+Unlike image data, which has spatial dependencies that aid the diffusion process, tabular data lacks this inherent structure, making the data generation process less efficient.
+Furthermore, diffusion models struggle with discrete categorical features, which are common in network traffic datasets, further hindering their performance [77].
+Gaussian Mixture Models also fail to capture the complex nature of network traffic data.
+These models assume that the data follows underlying Gaussian distributions, but they are unable to account for the multimodal and varied traffic behaviors typically found in network traffic datasets [78].
+For practical recommendations, techniques like SMOTE and ROS are still viable options for addressing class imbalance in network traffic datasets.
+When it comes to high-fidelity data generation, CTGAN and CopulaGAN are the top performers, surpassing other models in generating realistic and useful synthetic data.
+However, Bayesian Networks and Diffusion Models should be avoided for large-scale datasets due to their high computational costs and inefficiencies in handling such data.
+VII Future Work
 Building upon the insights derived from this study, potential directions for future investigations include:
@@
     Fidelity vs Privacy: Enhancing the fidelity and utility of synthetic data often poses privacy risks.
-Thus, it is imperative to identify an optimized balance between fidelity and privacy, particularly concerning cybersecurity datasets.
-This endeavor requires careful consideration to safeguard individual privacy while maintaining the data’s utility for cybersecurity purposes.
+Thus, it is imperative to identify an optimized balance between fidelity and privacy, particularly concerning cybersecurity network traffic datasets.
+This endeavor requires careful consideration to safeguard individual privacy while maintaining the data’s utility.
   4. 4.
@@
 This will allow a more critical understanding of how variables interact, particularly in complex datasets where correlations may not fully capture the underlying dependencies.
-IX Conclusion
-Our comparative study systematically explores the effectiveness of synthetic tabular data generation techniques within cybersecurity network traffic, covering both daily and malicious traffic scenarios.
-Through a detailed evaluation of several synthetic data generation methods across two significant datasets, we have revealed critical insights into their performances and the inherent trade-offs they entail.
-SMOTE and Cluster Centroids demonstrated substantial data similarity and class balance but suffer from the drawback of producing redundant, non-novel values.
-Conversely, CTGAN and TVAE outperformed the other methods, although they do not inherently ensure balanced class distribution.
-CastGAN emerges as a strong contender by maintaining internal feature dependencies, offering a more balanced approach.
-The insights from this study provide essential guidance for selecting synthetic data generation techniques based on specific cybersecurity requirements.
-This research significantly impacts the field by highlighting the strengths and limitations of each method, offering a clear framework for optimizing synthetic data generation in cybersecurity applications.
-The source code of this experiment is available at our https://github.com/AdenRajput/Comparative_Analysis.git.
-X Acknowledgements
+VIII Conclusion
+This study systematically evaluates synthetic network traffic data generation methods, comparing statistical and AI-based approaches.
+Our findings reveal that while statistical methods effectively handle class imbalance, they lack the ability to preserve complex feature dependencies.
+GAN-based models, particularly CTGAN and CopulaGAN, offer the best trade-off between fidelity and utility, making them preferable for generating realistic synthetic network traffic data.
+Diffusion models, despite their recent advances, exhibit significant computational challenges when applied to large-scale tabular datasets.
+The study highlights the need for hybrid approaches integrating statistical and AI techniques to improve class balance while maintaining data realism.
+Future research should explore optimizing generative AI loss functions, benchmarking real versus AI-generated class balance, and addressing privacy concerns while enhancing synthetic data fidelity.
+IX Acknowledgements
 The work presented here was partly financed by the European Celtic+ and Swedish Vinnova project ”CISSAN – Collective Intelligence Supported by Security Aware Nodes”.
+Appendix A Complete Graph Results for NSL-KDD Dataset
+Figure 5: Probability Distribution Comparison for NSL KDD (Part 1/18)
+Figure 6: Probability Distribution Comparison for NSL KDD (Part 2/18)
+Figure 7: Probability Distribution Comparison for NSL KDD (Part 3/18)
+Figure 8: Probability Distribution Comparison for NSL KDD (Part 4/18)
+Figure 9: Probability Distribution Comparison for NSL KDD (Part 5/18)
+Figure 10: Probability Distribution Comparison for NSL KDD (Part 6/18)
+Figure 11: Probability Distribution Comparison for NSL KDD (Part 7/18)
+Figure 12: Probability Distribution Comparison for NSL KDD (Part 8/18)
+Figure 13: Probability Distribution Comparison for NSL KDD (Part 9/18)
+Figure 14: Probability Distribution Comparison for NSL KDD (Part 10/18)
+Figure 15: Probability Distribution Comparison for NSL KDD (Part 11/18)
+Figure 16: Probability Distribution Comparison for NSL KDD (Part 12/18)
+Figure 17: Probability Distribution Comparison for NSL KDD (Part 13/18)
+Figure 18: Probability Distribution Comparison for NSL KDD (Part 14/18)
+Figure 19: Probability Distribution Comparison for NSL KDD (Part 15/18)
+Figure 20: Probability Distribution Comparison for NSL KDD (Part 16/18)
+Figure 21: Probability Distribution Comparison for NSL KDD (Part 17/18)
+Figure 22: Probability Distribution Comparison for NSL KDD (Part 18/18)
+Appendix B Complete Graph Results for CIC-IDS2017 Dataset
+Figure 23: Probability Distribution Comparison for CIC-IDS2017 (Part 1/7)
+Figure 24: Probability Distribution Comparison for CIC-IDS2017 (Part 2/7)
+Figure 25: Probability Distribution Comparison for CIC-IDS2017 (Part 3/7)
+Figure 26: Probability Distribution Comparison for CIC-IDS2017 (Part 4/7)
+Figure 27: Probability Distribution Comparison for CIC-IDS2017 (Part 5/7)
+Figure 28: Probability Distribution Comparison for CIC-IDS2017 (Part 6/7)
+Figure 29: Probability Distribution Comparison for CIC-IDS2017 (Part 7/7)
+Figure 30: Correlation Heatmap for CIC-IDS2017 (Part 1/2)
+Figure 31: Correlation Heatmap for CIC-IDS2017 (Part 2/2)